### v1

# Torque-based Graph Surgery: Enhancing Graph Neural Networks with Hierarchical Rewiring

Abstract
Graph Neural Networks (GNNs) have emerged as powerful tools for learning from graph-structured data, leveraging message passing to diffuse information and update node representations. However, most efforts have suggested that native interactions encoded in the graph may not be friendly for this process, motivating the development of graph rewiring methods. In this work, we propose a torque-driven hierarchical rewiring strategy, inspired by the notion of torque in classical mechanics, dynamically modulating message passing to improve representation learning in heterophilous graphs and enhance robustness against noisy graphs. Specifically, we define an interference-aware torque metric that integrates structural distance and energy scores to quantify the perturbation induced by edges, thereby encouraging each node to aggregate information from its nearest low-energy neighbors. We use the metric to hierarchically reconfigure each layer’s receptive field by judiciously pruning high-torque edges and adding low-torque links, suppressing propagation noise and boosting pertinent signals. Extensive evaluations on benchmark datasets show that our approach surpasses state-of-the-art methods on both heterophilous and homophilous graphs, and maintains high accuracy on noisy graph.
1 Introduction
Graph-structured data composed of vertices and edges encode entities and their relationships. Graph neural networks (GNNs) have emerged as a powerful framework for processing such data, with widespread applications in biomolecular modelling [1, 2], recommendation systems [3, 4] and beyond [5, 6, 7]. At the heart of GNNs lies message passing, which iteratively propagates and aggregates information along edges to enrich node representations. Therefore, the graph structure not only encodes entity interactions but also critically determines model performance [8, 9, 10].
In practice, however, raw graphs frequently harbour spurious or missing links arising from noise or sampling artefacts, compromising their effectiveness as substrates for message propagation. In response, recent work has devised diverse graph rewiring strategies that selectively remove and add edges to optimize message passing and boost predictive accuracy [11, 12, 13, 14]. Such dynamic topology adjustment is crucial not only for mitigating noise but also for addressing heterophily, where nodes with dissimilar labels or features tend to be connected [15, 16, 17]. In such scenarios, homophily-based GNNs can be misled by abundant heterophilous connections, yielding entangled representations and degraded classification accuracy.
To explore the factors that influence message propagation for guiding effective rewiring, we simulate attacks by injecting adversarial edges into raw graphs and visualize the resulting structural perturbations. As shown in Figure 1(a)-(b), adversarial edges in the homophilous Cora span node pairs with larger feature distances than native connections. A similar trend is observed in Wisconsin that inherently includes many heterophilous edges. To further uncover the attacker’s intention, we estimate node importance via the PageRank algorithm [18, 19] and derive an edge-level PageRank score by aggregating scores of each edge’s endpoints. As illustrated in Figure 1(c)-(d), adversarial edges exhibit two distinct patterns: in Cora, they preferentially connect high-PageRank nodes; whereas in Wisconsin, they appear predominantly among low-PageRank nodes, suggesting that PageRank cannot fully reveal the essence of adversarial edges. Moreover, real-world graphs inherently include spurious links even without external attacks, rendering PageRank insufficient to accurately assess the disruptive potential of edges. This motivates a key question: can we jointly leverage structural distance and a more expressive metric of edge perturbation to achieve fine-grained graph refinement?
Figure 1: Density distributions of distances and PageRank scores for normal vs. adversarial edges on Cora and Wisconsin, where edge PageRank is the sum of endpoint scores in Cora and the absolute difference in Wisconsin.
We draw inspiration from the concept of Torque in classical mechanics, mathematically defined as the cross product of a lever arm (the position vector from the axis of rotation to the point of force application) and a force. Recently, torque has found utility in biology [20, 21, 22] and spintronics [23, 24]. Heuristically, we extend this concept to graphs by treating the distance between nodes as the lever arm and the noise exerted by a neighboring node on a central node as the force. Their product yields a graph torque, which measures an edge’s disruptive intensity: higher torque flags greater interference. To our knowledge, this is the first work to integrate a physics-inspired torque into graph rewiring, enabling an interference-aware message passing.
Specifically, we devise a Torque-driven Hierarchical Rewiring strategy (THR) for GNNs (TorqueGNN), which dynamically refines message passing to excel in both heterophilous and noisy graph. In THR, each edge is endowed with two learnable attributes: its Euclidean distance as the “lever arm”, which accentuates long-range or heterophilous links, and an energy score as the “force”, which is computed by its endpoints reflecting high values to noisy or noise-sensitive nodes and low values to benign ones. Their product defines a torque to quantify the interference strength of edges, where larger torques flag less reliable connections. Leveraging this metric, TorqueGNN hierarchically reconfigures each layer’s receptive field via automatically removing undesirable edges that result in performance degrading and introducing low-torque significant connections, thus effecting interference-resistant and importance-aware propagation. This rewiring is performed end-to-end, where message passing operates on the continuously updated graph, while the evolving node representations enhance torque computation.
Contributions: 1) To our knowledge, we are the first effort to apply the torque in physics into graph rewiring, culminating in TorqueGNN that enhances GNNs resilience to both heterophily and noise. 2) We devise a hierarchical rewiring strategy that adaptively infers each layer’s receptive field through automatically pruning undesirable connections and adding significant edges. 3) Comprehensive experiments indicate that TorqueGNN outperforms advanced GNNs across standard benchmarks and maintains robustness in noisy graphs, with THR surpassing superior rewiring strategies.
2 Related Work
Standard message passing in GNNs, which aggregates information from local neighbourhoods, struggles to capture long-range dependencies. A common remedy is to stack multiple layers to expand the receptive field [25, 26, 27], but this approach frequently encounters fundamental limitations such as over-smoothing and over-squashing. To address these bottlenecks, graph rewiring techniques have recently emerged as a powerful strategy to restructure connectivity and enhance information flow. For example, Expander GNNs and ExPhormer perform graph rewiring by merging multi-hop neighbourhoods or injecting virtual nodes [28, 29, 30]. [31] adds edges based on spectral expansion to mitigate over-smoothing and over-squashing, while degree-preserving local edge-flip algorithms are developed [32]. [33, 34] analyze the root causes of over-squashing, demonstrating that both spatial and spectral rewiring can effectively counteract this bottleneck. Despite these advances, real-world graphs often remain inherently noisy, limiting the reliability of message passing.
Raw graph structures are frequently suboptimal substrates for message passing, as noisy edges and perturbed features can induce both spurious and missing links, thereby limiting the expressive power of GNNs [33, 35]. To mitigate the effect of such undesirable connections, [36] simultaneously perform graph rewiring and feature denoising to boost classification accuracy. Moreover, [36] highlights the challenge posed by heterophilous edges, wherein the aggregation of dissimilar node signals can lead to entangled representations and misclassifications. There are many methods specifically devised for heterophily. For example, [37, 38, 39] employ signed message propagation, assigning positive weights to homophilous links and negative weights to heterophilous ones. This enables differentiated updates on the heterophilous graphs, thereby amplifying similarity among homophilous nodes while suppressing similarity among heterophilous ones. However, [40] has shown that, although single-hop signed adjacency matrix aids in distinguishing features of different classes, the multi-hop propagation matrix introduced to expand the receptive field often degrades performance.
We are inspired by the torque in physics to design a new rewiring mechanism that hierarchically eliminates undesirable connections and incorporates task-relevant edges. By dynamically reshaping the receptive field during training, our method enhances the discriminative power of GNNs in noisy and heterophilous graphs.
3 Preliminaries
3.1 Notations
Let us define an undirected graph dataset as $\mathcal{G}=(V,\mathcal{E})$, comprising $N$ nodes $\{v_{i}\in V\}_{i=1}^{N}$ and $K$ edges $\{e_{k}\doteq\langle i,j\rangle\in{\mathcal{E}}\}_{k=1}^{K}$, where each edge $k$ encodes a connection between nodes $v_{i}$ and $v_{j}$. We denote the adjacency matrix by $\mathbf{A}\in\{0,1\}^{N\times N}$, where $A_{\langle i,j\rangle}=1$ iff nodes $v_{i}$ and $v_{j}$ are connected, 0 otherwise. Furthermore, $\widehat{\mathbf{A}}=\mathbf{A}+\mathbf{I}$ indicates $\mathbf{A}$ with added self-loops, and $\widetilde{\mathbf{A}}=\widehat{\mathbf{D}}^{-1/2}\widehat{\mathbf{A}}\widehat{\mathbf{D}}^{-1/2}$ denotes the symmetrically normalized adjacency matrix with $\widehat{{D}}_{\langle i,i\rangle}=\sum_{i=1}^{N}\widehat{A}_{\langle i,j\rangle}$. Each node is associated with a feature vector, and we write $\mathbf{X}\in\mathbb{R}^{N\times d}$ for the node feature matrix, where the $i$-th row $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the $d$-dimensional features of node $v_{i}$. Among $N$ nodes, $N_{lab}$ nodes are labeled, with ground-truth labels encoded in a matrix $\mathbf{Y}\in\mathbb{R}^{N_{lab}\times c}$, where each row $\mathbf{y}_{i}$ is a one-hot vector indicating the class label among $c$ categories.
3.2 Message Passing
Consider a graph with adjacency matrix $\mathbf{A}$ and node feature matrix $\mathbf{X}$. Message passing in a GNN proceeds by iteratively propagating and aggregating neighborhood information as
$\mathbf{h}_{i}^{(l+1)}=\text{Upd}\big{(}\mathbf{h}_{i}^{(l)},\sum_{v_{j}\in\mathcal{N}_{i}}\text{Agg}(\mathbf{h}_{j}^{(l)},A_{\langle i,j\rangle})\big{)},$ (1)
[…]
3.3 Energy-based Model
Energy-based model (EBM) [42] aims to establish an energy function $E(\mathbf{x},y):\mathbb{R}^{d}\rightarrow\mathbb{R}$, which maps an input instance $\mathbf{x}$ with a given label $y$ to a scalar value termed energy. Consequently, the Boltzmann distribution is expressed as
$p(y|\mathbf{x})=\frac{\text{exp}\big{(}-E(\mathbf{x},y)\big{)}}{\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}}=\frac{\text{exp}\big{(}-E(\mathbf{x},y)\big{)}}{\text{exp}\big{(}-E(\mathbf{x})\big{)}},$ (2)
where the denominator $\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}$ is the partition function. [43, 44] connect neural classifiers $f(\cdot)$, which map an input to a $c$-dimensional logit vector, with energy-based models by setting the energy as the negative logit value of the classifier $E(\mathbf{x},y)=-f(\mathbf{x})[y]$, where $f(\mathbf{x})[y]$ denotes the non-normalized logit corresponding to class $y$. Then, the energy of the input instance marginalizing over all possible labels becomes
$E(\mathbf{x})=-\text{log}\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}=-\text{log}\sum_{y^{\prime}}\text{exp}\big{(}f(\mathbf{x})[y^{\prime}]\big{)}.$ (3)
4 Methodology
In this section, we propose a novel graph rewiring strategy that unfolds in three key stages: (i) computing edge torques, (ii) rewiring propagation matrix, and (iii) adjusting message passing. The full algorithmic pseudocode is provided in Appendix A, while a detailed complexity analysis of TorqueGNN is included in Appendix B.
4.1 Derive Graph Torque
In classical mechanics, torque is defined as the vector cross product of a force and its lever arm:
$\mathbf{T}=\mathbf{r}\times\mathbf{F},\quad|\mathbf{T}|=|\mathbf{r}||\mathbf{F}|\sin\theta,$ (4)
where $\mathbf{r}$ denotes the lever arm vector connecting the point of application to the axis of rotation, $\mathbf{F}$ indicates the force vector, and $\theta$ is the angle between them. The magnitude of the torque directly governs an object’s tendency to rotate under the applied force. In GNNs, the systematic exploration of node interactions mirrors the lever arm–force relationship underlying torque in classical mechanics. Specifically, we can transpose torque to graphs by treating the distance $D_{\langle i,j\rangle}$ between a central node $v_{i}$ and its neighbor $v_{j}$ as a scalar “lever arm”, and the neighbor’s influence level $E_{\langle i,j\rangle}$ as a scalar “disturbing force” on $v_{i}$. In graphs, the notion of a disturbing force is abstract and lacks a well-defined direction, as message passing in GNNs does not involve physical "rotation" in the classical mechanical sense. Therefore, rather than computing the full vector torque via a cross product, we adopt a scalar approximation based on the product of distance and energy. Mathematically, for the edge $k$ connecting nodes $v_{i}$ and $v_{j}$, we define its torque value as follow,
$T_{e_{k}}=D_{\langle i,j\rangle}\times E_{\langle i,j\rangle},$ (5)
which measures the disturbance strength that the message carried by edge $e_{k}$ imposes on node $v_{i}$. Obviously, larger distances or higher disturbance both amplify torque, while edges maximizing both factors yield the greatest torque that represents the highest priority for graph rewiring. Crucially, the definition of $D$ and, in particular, $E$ in Eq. (5), remains a key challenge for accurate rewiring. Here, we employ an EBM to quantify node interference, and subsequently delineate the construction of both $D$ and $E$.
Metric 1: Distance. To mitigate the effect of noise in raw graphs and features, we estimate node distances using optimized representations. Specifically, for a central node $v_{i}$ and its neighbor $v_{j}$, we have the following formulation
$D_{\langle i,j\rangle}=\|\mathbf{h}_{i}-\mathbf{h}_{j}\|_{2},$ (6)
[…]
Metric 2: Energy. While graph-based message passing can markedly enhance node representations, it also risks amplifying structural and feature noise. To mitigate this, we posit nodes that persistently absorb noisy signals exhibit elevated interference strength relative to benign nodes. EBM, which assigns scalar energy values to inputs and has demonstrated remarkable robustness against adversarial perturbations [45, 46], provide a foundation for this purpose. Therefore, we introduce an EBM-derived energy score as another indicator of edge reliability, as shown in Eq. (3), which assigns low energy to benign samples and high energy to corrupted ones. In our framework, an energy-based GNN jointly performs node classification and energy estimation. For a network parameterized by $\mathbf{\Theta}$, the energy of node $v_{i}$ is defined as $E_{\mathbf{\Theta}}(\mathbf{x}_{i})=-\text{log}\sum_{y}\text{exp}(\mathbf{h}_{i}[y]\big{)}$. Here, $\mathbf{h}_{i}$ denotes the logit vector for node $v_{i}$. We then derive an edge-level energy metric by combining the energy scores of node pairs,
$E_{\langle i,j\rangle}=E_{\mathbf{\Theta}}(\mathbf{x}_{i})\times E_{\mathbf{\Theta}}(\mathbf{x}_{j}).$ (7)
4.2 Adjust Message Passing
Edge-removal High-order Rewiring. Herein, we propose an automated threshold learning mechanism that identifies the optimal number of edges to prune by pinpointing the largest successive torque gap. Specifically, we first rank all $K$ edges in descending order of their torque values to form a torque-sorted list (TSL), denoting its $k$-th entry entry as $\widetilde{e}_{k}$ with torque $\widetilde{T}_{e_{k}}$, so that $\widetilde{T}_{e_{1}}\geq\widetilde{T}_{e_{2}}\geq\cdots\geq\widetilde{T}_{e_{K}}$. We then calculate the torque gap between two consecutive links by
$G_{k,k+1}=\mu_{k}\times\frac{\widetilde{T}_{e_{k}}}{\widetilde{T}_{e_{k+1}}+\delta},$ (8)
where $\delta$ is a small constant to prevent division by zero, and the weight $\mu_{k}$ reflects the proportion of anomalous edges, those whose distance $D$, energy $E$ and torque $T$ all exceed their respective means, that are captured within the top $k$ torque-ranked set, emphasizing the boundary between desirable and undesirable connections. The computation formula of $\mu_{k}$ is defined as
$\begin{split}\mu_{k}&=\frac{|High\_e\cap Top\_k|}{|High\_e|},\\ \hfill High\_e&=\{e_{k}\doteq{\langle i,j\rangle}|D_{\langle i,j\rangle}\geq\bar{D},E_{\langle i,j\rangle}\geq\bar{E},T_{\langle i,j\rangle}\geq\bar{T},\langle i,j\rangle\in\mathcal{E}\},\\ Top\_k&=\{e_{k}\doteq{\langle i,j\rangle}|Top_{k}\{T_{\langle i,j\rangle}\},\langle i,j\rangle\in\mathcal{E}\},\end{split}$ (9)
where $\bar{D},\bar{E},\bar{T}$ denote the mean values of distance, energy and torque, respectively, computed over all $K$ edges. The set $High\_e$ comprises edges exhibiting above-average values across all three metrics, while $Top\_k$ contains the top $k$ connections in TSL. According to (8), we can identify the optimal cutoff by locating the largest torque gap $\mathcal{K}=\arg\max\limits_{0\leq k\leq K-1}G_{k,k+1}$, which separates the edge set into two groups: undesirable connections ($\widetilde{e}_{1},\cdots,\widetilde{e}_{\mathcal{K}}$) and desirable connections ($\widetilde{e}_{\mathcal{K}+1},\cdots,\widetilde{e}_{K}$).
In practice, multi-layer GNNs, such as APPNP [47], are widely adopted to enlarge the receptive field of graph convolutions. Building on the pairwise torque gap formulation introduced above, we extend this mechanism across multiple propagation layers. In specific, for each order $l$, we construct a dedicated propagation matrix that enables selective filtering of undesirable high-order interactions. Let $\mathcal{A}^{(l)}$ denote the refined propagation matrix used in $l$-th layer and $\mathcal{A}^{(0)}=\mathbf{A}$. We have current neighbors $v_{j}$ of a given node $v_{i}$ based on $\mathcal{A}^{(l)}$, and recompute two key properties as follows
$D_{\langle i,j\rangle}^{(l+1)}=\|\mathbf{h}_{i}^{(l)}-\mathbf{h}_{j}^{(l)}\|_{2},\quad E_{\langle i,j\rangle}^{(l+1)}=E_{\mathbf{\Theta}}^{(l)}(\mathbf{x}_{i})\times E_{\mathbf{\Theta}}^{(l)}(\mathbf{x}_{j}),$ (10)
[…]
$p_{kj}=\frac{\exp\left(\frac{\log\left(\pi_{kj}\right)+g_{kj}}{\tau}\right)}{\sum_{m=1}^{2}\exp\left(\frac{\log\left(\pi_{km}\right)+g_{km}}{\tau}\right)},\forall j=0,1,k\in\{1,2,\ldots,N\times t\},$ (11)
where $\tau$ is a temperature parameter controlling the sharpness of the Gumbel-Softmax distribution. $p_{k1}$ serves as a differentiable weight indicating the likelihood of selecting candidate edge $k$. Finally, we construct the rewired propagation matrix $\mathcal{A}^{(l+1)}$ by augmenting $\mathcal{A}^{(l+1)^{*}}$ with these probabilistically weighted candidate edges, followed by the standard renormalization procedure.
[…]
$\mathbf{h}_{i}^{(l+1)}$ $=\text{ReLU}\big{(}\sum\limits_{j\in\{v_{i}\}\cup\mathcal{N}_{i}^{(l+1)}}\alpha\mathbf{h}_{j}^{(l)}+(1-\alpha)\mathbf{h}_{i}^{(0)}\big{)}.$ (12)
The final node representations from the last layer are fed into a fully connected layer parameterized by $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ to yield the predicted class probabilities. These predictions are compared against the ground-truth labels using a cross-entropy loss minimized through gradient-based optimization.
4.3 Energy Calibration Objective
The energy function represents the unnormalized negative log-likelihood of node $v_{i}$, where a lower energy corresponds to a higher likelihood and, consequently, a greater probability of the sample being benign. In EBMs, the fundamental objective is to assign low energy to positive samples and high energy to others, thereby maximizing the likelihood of the observed data. This is typically achieved by minimizing the negative log-likelihood, formulated as
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\text{log}p_{\mathbf{\Theta}}(\mathbf{x}_{i}).$ (13)
[…]
$\begin{split}\mathcal{F}\big{(}p_{data}(\mathbf{x})\|p_{\mathbf{\Theta}}(\mathbf{x})\big{)}=\mathbb{E}_{p_{data}(\mathbf{x})}\Big{[}\frac{1}{2}\|\nabla_{\mathbf{\mathbf{x}}}\log p_{data}(\mathbf{x})-\nabla_{\mathbf{x}}\log p_{\mathbf{\Theta}}(\mathbf{x})\|^{2}\Big{]}.\end{split}$ (14)
To enlarge the energy signal of undesirable connections to facilitate more clearer distinction during edge selection, inspired by [46], we generate negative examples by randomly perturbing the graph structure, i.e., by adding or removing edges. Let $\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{i})$ denote the energy of node $v_{i}$ on this perturbed graph. We then define an energy-based score according to the following definition
Definition 1
(Energy Score): For a node $v_{i}$, its score $S_{\mathbf{\Theta}}(\mathbf{x}_{i})$ assigned by the energy model is surrogated by a gradient in the discrete space:
$S_{\mathbf{\Theta}}(\mathbf{x}_{i})=\nabla_{\mathbf{x}_{i}}E_{\mathbf{\Theta}}(\mathbf{x}_{i})=\Big{[}\frac{E_{\mathbf{\Theta}}(\mathbf{x}_{1})-\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{1})}{E_{\mathbf{\Theta}}(\mathbf{x}_{1})},\cdots,\frac{E_{\mathbf{\Theta}}(\mathbf{x}_{M})-\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{M})}{E_{\mathbf{\Theta}}(\mathbf{x}_{M})}\Big{]},$ (15)
where $M$ is the number of perturbed samples used. In fact, since $-\log p_{\mathbf{\Theta}}(\mathbf{x})=E_{\mathbf{\Theta}}(\mathbf{x})$, Eq. (15) equals to the following formula,
$\begin{split}S_{\mathbf{\Theta}}(\mathbf{x}_{i})=\nabla_{\mathbf{x}_{i}}E_{\mathbf{\Theta}}(\mathbf{x}_{i})=\Big{[}\frac{\log p_{\mathbf{\Theta}}(\mathbf{x}_{1})-\log\widehat{p}_{\mathbf{\Theta}}(\mathbf{x}_{1})}{\log p_{\mathbf{\Theta}}(\mathbf{x}_{1})},\cdots,\frac{\log p_{\mathbf{\Theta}}(\mathbf{x}_{M})-\log\widehat{p}_{\mathbf{\Theta}}(\mathbf{x}_{M})}{\log p_{\mathbf{\Theta}}(\mathbf{x}_{M})}\Big{]}.\end{split}$ (16)
Due to the energy score surrogate, the objective (14) of score matching is reformulated as follows
$\mathcal{F}\big{(}p_{data}(\mathbf{x}_{i})\|p_{\mathbf{\Theta}}(\mathbf{x}_{i})\big{)}=\mathbb{E}_{p_{data}(\mathbf{x}_{i})}\Big{[}\frac{1}{2}\|S_{data}(\mathbf{x}_{i})-S_{\mathbf{\Theta}}(\mathbf{x}_{i})\|^{2}\Big{]}.$ (17)
However, the real data is unavailable during training, and thus following [49, 46] and Definition 1, the final optimization formula is given by
$\mathcal{L}_{sm}=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\mathbf{x}_{i}}S_{\mathbf{\Theta}}(\mathbf{x}_{i})^{\top}\nabla_{\mathbf{x}_{i}}S_{\mathbf{\Theta}}(\mathbf{x}_{i})+\frac{1}{2}\|S_{\mathbf{\Theta}}(\mathbf{x}_{i})\|^{2}\Big{]}.$ (18)
By minimizing the surrogate loss (18), we not only approximate the original score matching objective of the EBM, but also explicitly amplify the energy signals associated with undesirable edges.
5 Experiments
Datasets. We evaluate our method on ten standard node classification benchmarks, including three homophilous graphs (Citeseer, Cora, and Pubmed) and seven heterophilous graphs (Texas, Wisconsin, Cornell, Actor, Tolokers, Questions, and Penn94). Among them, Tolokers, Questions, and Penn94 are large-scale datasets. Following previous work, we adopt the same data split strategy: 48% of the nodes are used for training, 32% for validation, and the remaining 20% for testing. The statistics of these datasets are summarized in Table 1, and further details can be found in Appendix C.2.
Table 1: Benchmark dataset statistics.
Datasets Edge Hom. #Nodes #Edges #Classes #Features
Texas 0.11 183 295 5 1,703
Wisconsin 0.21 251 466 5 1,703
Cornell 0.30 183 280 5 1,703
Actor 0.22 7,600 26,752 5 931
Tolokers 0.60 11,758 51,900 2 10
Questions 0.84 48,921 153,540 2 301
Penn94 0.47 41,554 1,362,229 2 4,814
Citeseer 0.74 3,327 4,676 7 3,703
Cora 0.81 2,708 5,278 6 1,433
Pubmed 0.80 19,717 44,327 3 500
Competitors. To evaluate the effectiveness of our model in capturing high-order interactions, we adopt APPNP as the downstream GNN, due to its suitability for deep graph architectures. We compare our approach against eleven baselines: a two-layer MLP; three classic homophily-oriented GNNs, including GCN [50], GAT [41] and APPNP [47]; and seven representative methods specifically designed for heterophilous graphs: H2GCN [51], GPR-GNN [52], FAGCN [39], ACM-GNN [38], OGNN [53], M2M-GNN [40], PCNet [54]. Furthermore, to assess the impact of our THR rewiring strategy, we compare it with other rewiring strategies: first-order spectral rewiring (FoSR) [55], joint denoising and rewiring (JDR) [36] and randomly removing edges (DropEdge) [56]. More details on all methods are provided in Appendix C.3.
Setups. We report node classification ACC, defined as the proportion of correctly predicted labels. For most benchmark datasets, models are trained using the Adam optimizer for 100 epochs with a learning rate of 0.001 and 512 hidden units. For larger datasets such as Tolokers, Questions, and Penn94, the number of training epochs is increased to 500 to ensure sufficient training. Detailed hyperparameter and environment configurations are provided in Appendix C.4, including the code link in C.1. Each experiment is repeated ten times, and we report the mean and standard deviation.
Node Classification Results. Table 2 reports test-set accuracy for three TorqueGNN variants, including edge-addition THR (A-THR), edge-removal THR (R-THR) and mixed THR (M-THR), alongside eleven state-of-the-art baselines across seven benchmark datasets. Several insights emerge: 1) On every dataset, at least one TorqueGNN variant appears among the top three performers, and TorqueGNN achieves the highest accuracy on the majority of benchmarks. 2) Three rewiring strategies yield substantial gains over APPNP and improvements are most pronounced on heterophilous graphs, underscoring the significance of our torque-based rewiring. 3) On Citeseer, Cora, and Pubmed, both A-THR and R-THR outperform the M-THR variant, suggesting that simultaneous addition and removal can inadvertently constrain information flow for citation networks. 4) In highly heterophilous datasets, although edge addition boosts APPNP, original undesirable links limit its effect. By contrast, R-THR achieves more high accuracy by judiciously deleting these connections. 5) In cases such as Pubmed and Cornell, R-THR attains the best results, likely because aggressive removal of noisy original edges without introducing additional candidates yields a cleaner propagation structure.
Table 2: Node classification results on benchmark datasets: Mean ACC % ± Standard Deviation %. he first-, second- and third-best accuracy are highlighted in red, green and blue, respectively.
Methods Citeseer Cora Pubmed Texas Wiscon. Cornell Actor
MLP 74.02±1.90 75.69±2.00 87.16±0.37 80.81±4.75 85.29±3.31 81.89±6.40 36.53±0.70
GCN 76.50±1.36 86.98±1.27 88.42±0.50 55.14±5.16 51.76±3.06 60.54±5.30 27.32±1.10
GAT 76.55±1.23 87.30±1.10 86.33±0.48 52.16±6.63 49.41±4.09 61.89±5.05 27.44±0.89
H2GCN 77.11±1.57 87.87±1.20 89.49±0.38 84.86±7.23 87.65±4.98 82.70±5.28 35.70±1.00
GPR-GNN 77.13±1.67 87.95±1.18 87.54±0.38 78.38±4.36 82.94±4.21 80.27±8.11 34.63±1.22
FAGCN 74.01±1.85 86.34±0.67 76.57±1.88 77.56±6.11 79.41±6.55 78.64±5.47 34.85±1.61
ACM-GCN 77.32±1.70 87.91±0.95 90.00±0.52 87.84±4.40 88.43±3.22 85.14±6.07 36.28±1.09
OGNN 77.31±1.73 88.37±0.75 90.15±0.38 86.22±4.12 88.04±3.63 87.03±4.73 37.99±1.00
M2M-GNN 77.20±1.80 88.12±1.00 90.35±0.60 89.19±4.50 89.01±4.10 86.48±6.10 36.72±1.60
PCNet 77.50±1.06 88.41±0.66 89.51±0.28 88.11±2.17 88.63±2.75 82.61±2.70 37.80±0.64
APPNP (backbone) 74.02±0.38 88.32±0.41 84.53±0.34 49.17±3.30 47.60±4.54 55.83±4.02 29.03±1.12
TorqueGNN w M-THR 75.86±0.23 84.78±0.28 87.52±0.06 91.11±1.67 89.60±1.20 89.44±1.67 36.72±0.16
TorqueGNN w R-THR 76.51±0.09 86.73±0.50 90.18±0.24 88.61±0.83 86.20±0.60 90.83±2.50 36.07±0.36
TorqueGNN w A-THR 78.03±0.05 88.47±0.28 89.38±0.02 86.11±0.01 82.00±0.01 83.61±3.15 36.24±0.04
Table 3: Node classification results on large-scale datasets: Mean ACC % (ROC AUC for imbalanced Questions) ± Standard Deviation %, where the optimal and suboptimal results are highlighted in red and green, respectively.
Penn94 Tolokers Questions
baseGNN 75.19 (0.17) 76.47 (1.16) 46.57 (0.71)
FoSR 75.21 (0.19) 76.64 (1.15) 46.95 (0.83)
JDR 75.19 (0.23) 78.06 (0.88) 46.98 (0.87)
DropEdge 78.46 (0.02) 77.03 (0.26) 60.01 (0.24)
THR 80.88 (0.01) 78.08 (0.16) 62.64 (0.89)
Results on Larger Graphs. Scalability of rewiring techniques on large graphs is crucial, particularly for end-to-end methods that dynamically add and remove edges during training. In TorqueGNN, the principal computational burden stems from computing high-order torques and corresponding gaps, which incurs a complexity of $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$ (see Appendix B for details). Despite this overhead, THR remains tractable for large graphs. Table 3 compares multiple rewiring schemes on larger heterophilous datasets: Both THR and DropEdge yield substantial accuracy improvements over the base GNN, yet THR consistently outperforms all alternatives. In contrast, FoSR and JDR achieve only marginal gains. Moreover, THR surpasses DropEdge, validating the effectiveness of the proposed torque-driven manner.
Figure 2: Results of TorqueGNN w/o THR and w THR on graphs with noisy edges.
Results on Adversarially Perturbed Graphs. Our torque-driven rewiring not only improves predictive accuracy but also enhances robustness to structural attacks by detecting and suppressing adversarial edges. To evaluate this, we employ MetaAttack [57] to inject adversarial links at a perturbation rate of 25% of the original edge count. Figure 2 compares classification accuracy under three settings: TorqueGNN without THR (w/o THR) on attacked graph, TorqueGNN with THR (w THR) on attacked graph and TorqueGNN w THR on normal graph. In all cases, we employ the mixed addition and deletion strategy for optimization. As shown, across both homophilous and heterophilous datasets, applying THR to the perturbed graph substantially improves classification performance, which demonstrates the effectiveness of our method in detecting and eliminating adversarial edges. Moreover, on heterophilous datasets, TorqueGNN w THR applied to the attacked graph even surpasses its performance on the normal graph, suggesting that the injected adversarial edges accentuate the distinction between desirable and undesirable connections, thereby enhancing the effectiveness of our rewiring mechanism.
Figure 3: Ablation study: Performance comparison of GCN and GAT with different rewiring schemes on four datasets.
Ablation Study. Table 2 has shown that all three THR variants markedly improve upon the APPNP backbone. To evaluate their broader applicability, we incorporate each strategy into two classic two-layer GNNs (GCN and GAT), and repeat the ablation studies in Figure 3. As illustrated, on the highly heterophilous Texas, Wisconsin, and Actor datasets, every rewiring variant substantially outperforms its base GNN. Even on the homophilous Pubmed, R-THR surpasses the base models, echoing the results in Table 2 and underscoring the broad applicability of our torque-driven optimization. We note, however, that A-THR and M-THR occasionally underperform relative to the baselines, likely because the current structure is already well-formed, and excessive edge additions may introduce noise, thereby degrading classification performance. Ablation study of significant components is included in Appendix C.5.
Parameter Analysis. To assess the influence of the key hyperparameter $\alpha$, we perform a sensitivity analysis shown in Figure 4. The figure examine the trade-off parameter $\alpha$, which balances the contribution of the learned hidden representation and the original input features, where a larger $\alpha$ increases the influence of the hidden representations. We observe that smaller graphs (e.g., Texas and Cornell), optimal accuracy is achieved at low $\alpha$ value ($0.01\leq\alpha\leq 0.05$), implying that raw node features provide sufficient discriminative power. In contrast, on medium-to-large graphs, better performance is observed when $\alpha\geq 0.5$, reflecting the necessity of high-order hidden representations to capture more complex community structures. Sensitivity analysis of other hyperparameters is presented in Appendix C.5.
[…]
6 Conclusion
In summary, we proposed a Torque-driven hierarchical rewiring graph neural network (TorqueGNN), which dynamically refined the graph structures to enhance representation learning on heterophilous and noisy graphs. By introducing an interference-aware torque metric, the product of the distance and the energy between node representations, TorqueGNN automatically removed undesirable connections and introduced beneficial ones during message passing. This hierarchical rewiring yielded interference-resilient, importance-aware propagation tailored to each layer’s receptive field. Extensive evaluations across homophilous and heterophilous benchmark datasets demonstrated that TorqueGNN consistently outperformed state-of-the-art baselines and maintained robust performance on adversarial graphs. We state the limitations of our model in Appendix D.
Appendix A Algorithm
Algorithm 1 outlines the complete workflow of TorqueGNN.
Input: Node features $\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$, candidate edge set $\mathcal{T}$, ground truth matrix $\mathbf{Y}$, hyperparameters $L$ and $\alpha$.
Output: The predicted class label.
1 Initialize network parameters $\mathbf{\Theta},\mathbf{\Phi}$;
2 $\mathbf{h}_{i}^{(0)}=\text{ReLU}(\mathbf{x}_{i}\mathbf{\Theta})$;
3 for $l=1\rightarrow L$ do
[…]
5   Compute pairwise distance $D_{\langle i,j\rangle}^{(l)}$ and energy $E_{\langle i,j\rangle}^{(l)}$ with Eqs. (6) and (7);
6   Compute the $l$th order torques with Eq. (5) and sort them. // Torque computation
7   Gain the largest torque gap $\mathcal{K}$ with Eqs. (8)-(9);
8   Remove the top $\mathcal{K}$ edges to gain $\mathcal{A}^{(l)^{*}}$. // Removing undesirable edges
9   Compute the sampling probability of candidate edges with Eq. (11);
10   Add beneficial candidate edges to form the refined propagation matrix $\mathcal{A}^{(l)}$. // Adding desirable connections
11   Update node representation $\mathbf{h}_{i}^{(l)}$ with Eq. (12). // Message passing
[…]
14 for $t=1\rightarrow T$ do
[…]
16
17Obtain $\mathbf{\hat{y}}_{i}=\text{Softmax}\left(\mathbf{h}^{(L)}_{i}\mathbf{\Phi}\right)$;
return The predicted class label of the $i$-th node is given by $\arg\max{\hat{\mathbf{y}}_{i}}$.
Algorithm 1 TorqueGNN
Appendix B Complexity Analysis
The dominant computational cost of TorqueGNN lies in: 1) Torque computation and graph rewiring. For each order $l$, we compute torque values only on the edges in $\mathcal{A}^{(l)}$, costing $\mathcal{O}(|\mathcal{A}^{(l)}|)$, and then sort these values in $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$. When adding edges, if the candidate set size is $B$, the combined probability calculation and sorting cost is $\mathcal{O}(B+B\log B)$. 2) Message passing on the rewired graph $\mathcal{A}^{(l)}$. Initial representations $\mathbf{H}^{(0)}$ are obtained by a fully connected layer with parameter $\mathbf{\Theta}\in\mathbb{R}^{d\times m}$ on $\mathbf{X}\in\mathbb{R}^{N\times d}$, at cost $\mathcal{O}(Ndm)$. Aggregation over $\mathcal{A}^{(l)}$ then costs $\mathcal{O}(m|\mathcal{A}^{(l)}|)$ per layer. The output layer with $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ requires $\mathcal{O}(Nmc)$. Putting these together for an $L$-layer network and assuming $B\ll|\mathbf{A}^{l}|$ for all $l$, the overall complexity is $\mathcal{O}(LNdm+\sum_{l=1}^{L}|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$.
Appendix C More Experimental Results
C.1 Configures
We construct a series of experiments to assess the proposed TorqueGNN. Our model is implemented in PyTorch on a workstation with AMD Ryzen 9 5900X CPU (3.70GHz), 64GB RAM and RTX 3090GPU (24GB caches). Our code is available at https://anonymous.4open.science/r/TorqueGNN-F60C/README.md.
C.2 Datasets
  • Homophilous Datasets. Citeseer, Cora and Pubmed are three citation networks, and they are published in [58]. Specifically,
    • Citeseer comprises 3,327 publications classified into six categories, with each paper encoded by a 3,703-dimensional binary word-presence vector.
    • Cora consists of 2,708 scientific publications classified into seven research topics. Each paper is represented by a 1,433-dimensional binary feature vector indicating the presence of specific word
    • Pubmed is a larger citation network of 19,717 diabetes-related articles labeled among three classes. Papers are described by 500-dimensional term frequency–inverse document frequency feature vectors, and citation edges capture scholarly references.
  • Heterophilous Datasets
    • Texas, Wisconsin, Cornell are WebKB datasets used in [59], where nodes correspond to individual web pages and edges correspond to the hyperlinks between them. Every node is described by a bag-of-words feature vector extracted from its page content, and each page has been manually labeled into one of five categories.
    • Actor [60] is the actor-only induced subgraph of a film–director–actor–writer network on Wikipedia, where each node represents an actor and an undirected edge connects two actors if they co-occur on the same Wikipedia page.
    • Tolokers [61] is built from the Toloka crowdsourcing platform, comprising 11,758 nodes and 519,000 edges that link workers who collaborated on the same task. Each node carries a 10-dimensional feature vector and is assigned one of two labels based on whether the worker was banned.
    • Questions [61] is an interaction graph of users on the Yandex Q question-answering platform, comprising 48,921 nodes and 153,540 edges that link users who interacted on the same question. Each node carries a 301-dimensional feature vector and a binary label for node classification.
    • Penn94 [62] is a subgraph of the Facebook100 dataset featuring 41,554 university students as nodes, connected by 1,362,229 undirected friendship edges. Each node is described by a five-dimensional feature vector and labeled by the gender of the students.
C.3 Methods
C.3.1 GNNs for Homophilous and Heterophilous Graphs
GCN generalize convolutional neural networks to graph-structured data by iteratively aggregating feature information from each node’s local neighborhood,
$\mathbf{h}^{(l)}_{i}=\sigma(\widetilde{\mathbf{A}}\mathbf{h}^{(l-1)}_{i}\mathbf{W}^{(l)}),$ (19)
where $\mathbf{W}$ is the learnable parameter matrix.
GAT leverages attention mechanisms to weight the influence of neighboring nodes during message passing. Its core rule is:
$\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}^{(l)}\right),\alpha_{ij}=\frac{\exp\left(\operatorname{LeakyReLU}\left(a^{\top}\left[\mathbf{W}\mathbf{h}_{i}^{(l)}\|\mathbf{W}\mathbf{h}_{j}^{(l)}\right]\right)\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\operatorname{LeakyReLU}\left(a^{\top}\left[\mathbf{W}\mathbf{h}_{i}^{(l)}\|\mathbf{W}\mathbf{h}_{k}^{(l)}\right]\right)\right)}.$ (20)
APPNP first achieves the feature transformation by:
$\mathbf{H}^{(0)}=\mathbf{X}\mathbf{W},$ (21)
and then propagating message via a Personalized PageRank scheme:
$\mathbf{H}^{(l)}=(1-\alpha)\mathbf{P}\mathbf{H}^{(l-1)}+\alpha\mathbf{H}^{(0)}.$ (22)
Here, $\mathbf{P}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ is the symmetrically normalized adjacency matrix and $\alpha$ is a trade-off hyperparameter.
H2GCN contains three designs, and the first design aims to separate ego- and neighbor-message:
$\mathbf{h}_{i}^{(l)}=[\mathbf{h}_{i}^{(l-1)}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}_{i}\})]$ (23)
Second, H2GCN directly aggregates neighbors of different hops, which enhances high-order information:
$\mathbf{h}_{i}^{(l)}=[\mathbf{h}_{i}^{(l-1)}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}^{1}_{i}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}^{2}_{i}||\cdots\})].$ (24)
Finally, the output is gained by combining the intermediate representations of all layers:
$\mathbf{h}_{i}=[\mathbf{h}_{i}^{(1)}||\mathbf{h}_{i}^{(2)}||\cdots||\mathbf{h}_{i}^{(L)}].$ (25)
GPR-GNN generalizes personalized PageRank by treating each hop’s contribution as a learnable parameter:
$\mathbf{H}=\sum_{l=1}^{L}\gamma^{l}\mathbf{P}\mathbf{H}^{(0)},\mathbf{H}^{(0)}=\mathbf{H}\mathbf{W},$ (26)
where $\gamma^{l}\mathbf{P}$ measures the propagation coefficient for the connection between nodes $v_{i}$ and $v_{j}$.
FAGCN adaptively balances low- and high-frequency graph signals via a learnable mixing coefficient:
$\mathbf{h}_{i}^{(l)}=\eta\mathbf{h}_{i}^{(0)}+\sum_{j\in\mathcal{N}_{i}}\frac{\alpha_{ij}}{\sqrt{d_{i}d_{j}}}\eta\mathbf{h}_{i}^{(l-1)},$ (27)
where $\alpha_{ij}\in(-1,1)$ is a learnable coefficient.
ACM-GNN dynamically learns how to weight multiple propagation channels:
$\mathbf{H}^{(l)}=\operatorname{ReLU}\left(\operatorname{diag}\left(\alpha_{H}^{(l)}\right)(\mathbf{I}-\mathbf{P})\mathbf{H}^{(l-1)}\mathbf{W}_{H}^{(l)}+\operatorname{diag}\left(\alpha_{L}^{(l)}\right)\mathbf{P}\mathbf{H}^{(l-1)}\mathbf{W}_{L}^{(l)}+\operatorname{diag}\left(\alpha_{I}^{(l)}\right)\mathbf{H}^{(l-1)}\mathbf{W}_{I}^{(l)}\right).$ (28)
Here, $\alpha$ is adaptive mixing weight.
OGNN introduces an order-aware message passing by sorting neighbors according to learned importance scores before aggregation:
$\mathbf{h}_{i}=\mathbf{h}_{i}^{(1)}\odot g_{i}^{(1)}+\mathbf{h}_{i}^{(2)}\odot\left(g_{i}^{(2)}-g_{i}^{(1)}\right)\ldots+\mathbf{h}_{i}^{(L)}\odot\left(g_{i}^{(KL}-g_{i}^{(L-1)}\right),$ (29)
where $g_{i}^{(l)}$ is a vector with a split point $p_{i}^{(l)}$ so that its first entries are 0 and its remaining entries are 1.
M2MGNN partitions neighbor embeddings into class-aligned “chunks” via learnable soft labels:
$\mathbf{h}_{i}^{(l)}=\operatorname{ReLU}\left((1-\beta)\mathbf{h}_{i}^{(0)}+\beta\mathbf{m}_{i}^{(l)}\right),\mathbf{C}_{it}^{(l)}=\sum_{j\in\mathcal{N}_{i}}\mathbf{s}_{t}^{(l)}(i,j)\hat{\mathbf{h}}_{j}^{(k)},\quad\mathbf{m}_{i}^{(l)}=\|_{t=1}^{C}\mathbf{C}_{it}^{(l)},$ (30)
where $\mathbf{s}_{t}^{(l)}(i,j)$ is the score of node $v_{i}$ towards $v_{j}$ in the $t$-th chunk.
PCNet introduces a unified two-fold filtering framework that simultaneously performs homophilous low-pass smoothing and heterophilous high-pass aggregation:
$\mathbf{H}=(\sum_{k=1}^{K}\theta_{k}\mathrm{PC}_{k}(\tilde{L})+\theta_{\mathrm{0}})\mathbf{X}\mathbf{W}.$ (31)
Here, $\mathrm{PC}_{k}(\tilde{L})$ denotes the $n$-th Poisson–Charlier polynomial of the normalized Laplacian $\tilde{L}$, and $\{\theta_{k}\}_{k=0}^{K}$ are learnable filter coefficients.
C.3.2 Rewiring Strategies
FoSR develope a spectral-expansion–based rewiring algorithm that added edges to alleviate over-squashing, and integrate it with a relational GNN to provably avoid oversmoothing.
JDR simultaneously refine graph topology and node features by aligning the principal eigenspaces of their matrices, boosting GNNs across both homophilous and heterophilous networks.
DropEdge randomly remove edges at each training epoch to act as both data augmentation and message-passing reduction, which is used to mitigate over-fitting and over-smooting problems.
C.4 Hyperparameters
In the subsection, we list the detailed hyperparameters used for the experiments and they are also provided in code. The hyperparameters can be found in Table 4.
Table 4: Hyperparameters of TorqueGNN in 10 datasets.
Datasets Lr Wd Dropout $L$ $t$ $r$ epochs
Texas 0.001 5e-4 0.7 8 5 0.2 100
Wisconsin 0.001 0.5 0.5 8 5 0.2 100
Cornell 0.001 0.05 0.7 8 2 0.2 100
Actor 0.001 0.05 0.1 8 5 0.2 100
Citeseer 0.001 5e-8 0.1 8 2 0.2 100
Cora 0.001 0.1 0.1 8 2 0.2 100
Pubmed 0.001 0.08 0.1 12 5 0.2 100
Tolokers 0.001 5e-8 0.1 2 2 0.2 500
Questions 0.001 5e-8 0.1 2 2 0.2 500
Penn94 0.001 5e-8 0.1 2 2 0.2 500
C.5 Experiments
Ablation Study. To validate the effectiveness of our proposed torque metric, we conduct ablation studies comparing three edge-scoring schemes. As reported in Table 5, reliance on structural distance alone yields the poorest classification performance on most benchmarks, reflecting the fact that certain long-range connections convey meaningful semantics. Introducing the energy metric markedly improves accuracy, confirming its significance as a measure of edge perturbation strength. Moreover, on Tolokers and Pubmed, the energy-only variant achieves the highest accuracy, slightly surpassing the combined torque score, likely because some low-interference connections carrying critical semantic information are inadvertently pruned when the distance is too large. Crucially, our torque score, integrating both distance and energy, consistently outperforms either criterion besides Tolokers and Pubmed, thereby validating its superior ability to accurately identify disruptive edges, guide interference-aware rewiring, and enhance overall classification accuracy.
Table 5: Ablation Study: Performance of TorqueGNN equipped with different edge-scoring metrics; the best results are shown in bold.
Datasets Distance Energy Torque (Distance $\times$ Energy)
Texas 88.61 (0.70) 89.16 (0.69) 91.11 (1.67)
Wisconsin 79.20 (2.40) 84.60 (1.80) 89.60 (1.20)
Cornell 81.94 (4.17) 83.05 (0.83) 89.44 (1.67)
Actor 36.69 (0.40) 35.48 (0.07) 36.72 (0.16)
Tolokers 72.46 (0.04) 76.11 (0.28) 75.28 (0.02)
Penn94 81.11 (0.09) 81.20 (0.15) 81.30 (0.19)
Questions 63.11 (1.26) 62.65 (0.53) 63.54 (0.85)
Citeseer 73.35 (0.24) 76.53 (0.42) 76.84 (0.60)
Cora 78.30 (0.59) 84.50 (0.00) 85.39 (0.30)
Citeseer 73.35 (0.24) 76.53 (0.42) 76.84 (0.60)
Pubmed 88.01 (0.11) 88.52 (0.16) 88.36 (0.08)
Parameter Sensitivity. Figure 5 explores the effect of network depth $L$. Across most datasets, performance peaks at $L=8$, thereby validating the effectiveness of our THR strategy for mining high-order information. Beyond eight layers, accuracy declines, primarily due to training challenges typical of deep networks, such as over-smoothing. While our rewiring mechanism mitigates over-smoothing to some extent, it is not explicitly engineered to resolve this phenomenon; overcoming depth-related bottlenecks remains a direction for future work.
Figure 5: Parameter sensitivity: Performance curves on five datasets with layers changing in $2\}$ , 4 , 8 , 16 , 32 }.
Appendix D Broader Impact Statement and Limitations
This study aims to enhance message passing in graph neural networks through graph rewiring. As a result, it contributes to better performance and broader applicability of GNNs across a wide range of tasks, including recommendation systems, molecular property prediction, traffic forecasting, and social network.
THR depends on sorting operations that impose considerable computational and memory burdens on large, densely connected graphs. Additionally, by assigning equal weight to distance and energy metrics, THR may inadvertently prune edges that carry essential semantic information and exhibit low interference when the distance is too large. How to optimally balance these two metrics in accordance with the characteristics of different datasets remains an open challenge.

### v3

@@
-Torque-based Graph Surgery: Enhancing Graph Neural Networks with Hierarchical Rewiring
+GraphTorque:: Torque-Driven Rewiring Graph Neural Network
 Abstract
@@
 However, most efforts have suggested that native interactions encoded in the graph may not be friendly for this process, motivating the development of graph rewiring methods.
-In this work, we propose a torque-driven hierarchical rewiring strategy, inspired by the notion of torque in classical mechanics, dynamically modulating message passing to improve representation learning in heterophilous graphs and enhance robustness against noisy graphs.
-Specifically, we define an interference-aware torque metric that integrates structural distance and energy scores to quantify the perturbation induced by edges, thereby encouraging each node to aggregate information from its nearest low-energy neighbors.
-We use the metric to hierarchically reconfigure each layer’s receptive field by judiciously pruning high-torque edges and adding low-torque links, suppressing propagation noise and boosting pertinent signals.
-Extensive evaluations on benchmark datasets show that our approach surpasses state-of-the-art methods on both heterophilous and homophilous graphs, and maintains high accuracy on noisy graph.
+In this work, we propose a torque-driven hierarchical rewiring strategy, inspired by the notion of torque in classical mechanics, dynamically modulating message passing to improve representation learning in heterophilous and homophilous graphs.
+Specifically, we define the torque by treating the feature distance as a “lever arm vector” and the neighbor feature as a “force vector” weighted by the homophily disparity between nodes.
+We use the metric to hierarchically reconfigure each layer’s receptive field by judiciously pruning high-torque edges and adding low-torque links, suppressing the impact of irrelevant information and boosting pertinent signals during message passing.
+Extensive evaluations on benchmark datasets show that the proposed approach surpasses state-of-the-art rewiring methods on both heterophilous and homophilous graphs.
 1 Introduction
@@
 In practice, however, raw graphs frequently harbour spurious or missing links arising from noise or sampling artefacts, compromising their effectiveness as substrates for message propagation.
-In response, recent work has devised diverse graph rewiring strategies that selectively remove and add edges to optimize message passing and boost predictive accuracy [11, 12, 13, 14].
-Such dynamic topology adjustment is crucial not only for mitigating noise but also for addressing heterophily, where nodes with dissimilar labels or features tend to be connected [15, 16, 17].
-In such scenarios, homophily-based GNNs can be misled by abundant heterophilous connections, yielding entangled representations and degraded classification accuracy.
-To explore the factors that influence message propagation for guiding effective rewiring, we simulate attacks by injecting adversarial edges into raw graphs and visualize the resulting structural perturbations.
-As shown in Figure 1(a)-(b), adversarial edges in the homophilous Cora span node pairs with larger feature distances than native connections.
-A similar trend is observed in Wisconsin that inherently includes many heterophilous edges.
-To further uncover the attacker’s intention, we estimate node importance via the PageRank algorithm [18, 19] and derive an edge-level PageRank score by aggregating scores of each edge’s endpoints.
-As illustrated in Figure 1(c)-(d), adversarial edges exhibit two distinct patterns: in Cora, they preferentially connect high-PageRank nodes; whereas in Wisconsin, they appear predominantly among low-PageRank nodes, suggesting that PageRank cannot fully reveal the essence of adversarial edges.
-Moreover, real-world graphs inherently include spurious links even without external attacks, rendering PageRank insufficient to accurately assess the disruptive potential of edges.
-This motivates a key question: can we jointly leverage structural distance and a more expressive metric of edge perturbation to achieve fine-grained graph refinement?
-Figure 1: Density distributions of distances and PageRank scores for normal vs.
-adversarial edges on Cora and Wisconsin, where edge PageRank is the sum of endpoint scores in Cora and the absolute difference in Wisconsin.
-We draw inspiration from the concept of Torque in classical mechanics, mathematically defined as the cross product of a lever arm (the position vector from the axis of rotation to the point of force application) and a force.
-Recently, torque has found utility in biology [20, 21, 22] and spintronics [23, 24].
-Heuristically, we extend this concept to graphs by treating the distance between nodes as the lever arm and the noise exerted by a neighboring node on a central node as the force.
-Their product yields a graph torque, which measures an edge’s disruptive intensity: higher torque flags greater interference.
+In response, recent work has devised diverse graph rewiring strategies that selectively remove and add edges to optimize message passing and boost predictive accuracy [11, 12, 13, 14, 15].
+Such dynamic topology adjustment is crucial not only for mitigating spurious connections but also for addressing heterophily, where nodes with dissimilar labels or features tend to be connected [16, 17, 18].
+In such scenarios, homophily‑based GNNs can be misled by abundant heterophilous connections, yielding entangled representations and degraded classification accuracy.
+One of the core challenges in graph rewiring is quantifying the impact of edges on message passing.
+A key factor in this process is the similarity between node pairs, often measured using the Euclidean distance, a commonly used metric for assessing similarity.
+In general, the larger the distance between nodes, the weaker their interaction strength, and the less useful information can be transmitted, as supported by previous studies that employed node similarity as a proxy for edge weights [19, 20].
+To intuitively observe this, we simulate adversarial attacks by injecting adversarial edges into raw graphs and visualize the distance distribution of the edges, enabling us to examine whether adversarial and original edges exhibit distinct distributional patterns.
+As shown in Fig.
+1(a)–(d), the distribution trends in both homophilous datasets (Cora and PubMed) and heterophilous datasets ( Wisconsin and Texas ) consistently indicate that adversarial edges (in red) tend to connect node pairs with larger feature distances.
+These observations suggest that adversarial attacks preferentially create long-range links so that they disrupt message passing at their target nodes.
+Furthermore, we observe that normal edges in heterophilous datasets also exhibit a distribution skewed toward larger distances, more pronounced than in homophilous datasets.
+This is because heterophilous graphs contain a much higher proportion of heterophilous edges, which typically span node pairs with low similarity (i.e., large distances).
+Given that a minority of long-range neighbors can convey crucial information while nearby neighbors may propagate misleading signals, the feature quality of neighboring nodes should be another key factor in assessing edge significance.
+Figure 1: Density distributions of distances for normal vs. adversarial edges on homophilous graphs (a) Cora and (b) Pubmed and heterophilous graphs (c) Wisconsin and (d) Texas.
+This brings to mind the concept of Torque in classical mechanics, which is mathematically defined as the cross product of a lever arm (the position vector from the axis of rotation to the point of force application) and a force.
+Recently, torque has found applications in fields such as biology [21, 22, 23] and spintronics [24, 25].
+Heuristically, we extend this concept to graphs by treating the distance vector between nodes as the lever arm and the feature vector of a neighboring node as the force.
+Their product yields a graph torque, which measures an edge’s negative impact: higher torque flags greater interference.
 To our knowledge, this is the first work to integrate a physics-inspired torque into graph rewiring, enabling an interference-aware message passing.
-Specifically, we devise a Torque-driven Hierarchical Rewiring strategy (THR) for GNNs (TorqueGNN), which dynamically refines message passing to excel in both heterophilous and noisy graph.
-In THR, each edge is endowed with two learnable attributes: its Euclidean distance as the “lever arm”, which accentuates long-range or heterophilous links, and an energy score as the “force”, which is computed by its endpoints reflecting high values to noisy or noise-sensitive nodes and low values to benign ones.
-Their product defines a torque to quantify the interference strength of edges, where larger torques flag less reliable connections.
-Leveraging this metric, TorqueGNN hierarchically reconfigures each layer’s receptive field via automatically removing undesirable edges that result in performance degrading and introducing low-torque significant connections, thus effecting interference-resistant and importance-aware propagation.
+Specifically, we devise a Torque‑driven Hierarchical Rewiring strategy (THR) for GNNs, which dynamically refines message passing to excel in both homophilous and heterophilous graph.
+In THR, each edge is assigned a torque value that quantifies its interference strength, with larger torques indicating less reliable connections.
+Sepcifically, we define torque by treating the difference between node representations as a “lever arm vector”, which emphasizes long-range or heterophilous links.
+Meanwhile, the neighbor feature is treated as a “force vector” weighted by the disparity in the homophily ratios between node pairs.
+This disparity captures the difference in their local label homophily, which has been theoretically shown to jointly influence the expressive power of GNNs together with feature distance.
+Leveraging this torque value, THR hierarchically reconfigures each layer’s receptive field via automatically removing undesirable edges that degrade performance and introducing low-torque significant connections, thus effecting interference-resistant and importance-aware propagation.
 This rewiring is performed end-to-end, where message passing operates on the continuously updated graph, while the evolving node representations enhance torque computation.
-Contributions: 1) To our knowledge, we are the first effort to apply the torque in physics into graph rewiring, culminating in TorqueGNN that enhances GNNs resilience to both heterophily and noise.
-2) We devise a hierarchical rewiring strategy that adaptively infers each layer’s receptive field through automatically pruning undesirable connections and adding significant edges.
-3) Comprehensive experiments indicate that TorqueGNN outperforms advanced GNNs across standard benchmarks and maintains robustness in noisy graphs, with THR surpassing superior rewiring strategies.
+Contributions: 1) To the best of our knowledge, we are the first to apply the concept of torque from physics to graph rewiring, resulting in THR, which enhances GNNs’ resilience to both homophily and heterophily.
+2) We propose a hierarchical rewiring strategy that adaptively determines each layer’s receptive field by automatically pruning undesirable connections and adding significant edges.
+3) Comprehensive experiments indicate that THR improves the performance of various GNNs and outperforms existing state-of-the-art rewiring strategies.
 2 Related Work
 Standard message passing in GNNs, which aggregates information from local neighbourhoods, struggles to capture long-range dependencies.
-A common remedy is to stack multiple layers to expand the receptive field [25, 26, 27], but this approach frequently encounters fundamental limitations such as over-smoothing and over-squashing.
+A common remedy is to stack multiple layers to expand the receptive field [26, 27, 28], but this approach frequently encounters fundamental limitations such as over‑smoothing and over‑squashing.
 To address these bottlenecks, graph rewiring techniques have recently emerged as a powerful strategy to restructure connectivity and enhance information flow.
-For example, Expander GNNs and ExPhormer perform graph rewiring by merging multi-hop neighbourhoods or injecting virtual nodes [28, 29, 30].
+For example, Expander GNNs and ExPhormer perform graph rewiring by merging multi‑hop neighbourhoods or injecting virtual nodes [29, 30, 31].
 [32] adds edges based on spectral expansion to mitigate over-smoothing and over-squashing, while degree-preserving local edge-flip algorithms are developed [33].
 [34, 35] analyze the root causes of over-squashing, demonstrating that both spatial and spectral rewiring can effectively counteract this bottleneck.
-Despite these advances, real-world graphs often remain inherently noisy, limiting the reliability of message passing.
-Raw graph structures are frequently suboptimal substrates for message passing, as noisy edges and perturbed features can induce both spurious and missing links, thereby limiting the expressive power of GNNs [33, 35].
-To mitigate the effect of such undesirable connections, [36] simultaneously perform graph rewiring and feature denoising to boost classification accuracy.
 Moreover, [36] highlights the challenge posed by heterophilous edges, wherein the aggregation of dissimilar node signals can lead to entangled representations and misclassifications.
-There are many methods specifically devised for heterophily.
-For example, [37, 38, 39] employ signed message propagation, assigning positive weights to homophilous links and negative weights to heterophilous ones.
+To mitigate the effect of such undesirable connections, [14] compares the neighborhood feature distribution and neighborhood label distribution between node pairs to pruning heterophilous edges and adding homophilous edges.
+[37, 38, 39] employ signed message propagation, assigning positive weights to homophilous links and negative weights to heterophilous ones.
 This enables differentiated updates on the heterophilous graphs, thereby amplifying similarity among homophilous nodes while suppressing similarity among heterophilous ones.
-However, [40] has shown that, although single-hop signed adjacency matrix aids in distinguishing features of different classes, the multi-hop propagation matrix introduced to expand the receptive field often degrades performance.
-We are inspired by the torque in physics to design a new rewiring mechanism that hierarchically eliminates undesirable connections and incorporates task-relevant edges.
-By dynamically reshaping the receptive field during training, our method enhances the discriminative power of GNNs in noisy and heterophilous graphs.
+However, [40] has shown that, although single‑hop signed adjacency matrix aids in distinguishing features of different classes, the multi‐hop propagation matrix introduced to expand the receptive field often degrades performance.
+We are inspired by the torque in physics to design a new rewiring mechanism that hierarchically eliminates undesirable connections and incorporates task‑relevant edges.
+By dynamically reshaping the receptive field during training, our method enhances the discriminative power of GNNs on both homophilous and heterophilous graphs.
 3 Preliminaries
@@
 $\mathbf{h}_{i}^{(l+1)}=\text{Upd}\big(\mathbf{h}_{i}^{(l)},\sum_{v_{j}\in\mathcal{N}_{i}}\text{Agg}(\mathbf{h}_{j}^{(l)},A_{\langle i,j\rangle})\big),$ (1)
-[…]
-3.3 Energy-based Model
-Energy-based model (EBM) [42] aims to establish an energy function $E(\mathbf{x},y):\mathbb{R}^{d}\rightarrow\mathbb{R}$, which maps an input instance $\mathbf{x}$ with a given label $y$ to a scalar value termed energy.
-Consequently, the Boltzmann distribution is expressed as
-$p(y|\mathbf{x})=\frac{\text{exp}\big{(}-E(\mathbf{x},y)\big{)}}{\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}}=\frac{\text{exp}\big{(}-E(\mathbf{x},y)\big{)}}{\text{exp}\big{(}-E(\mathbf{x})\big{)}},$ (2)
-where the denominator $\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}$ is the partition function.
-[43, 44] connect neural classifiers $f(\cdot)$, which map an input to a $c$-dimensional logit vector, with energy-based models by setting the energy as the negative logit value of the classifier $E(\mathbf{x},y)=-f(\mathbf{x})[y]$, where $f(\mathbf{x})[y]$ denotes the non-normalized logit corresponding to class $y$.
-Then, the energy of the input instance marginalizing over all possible labels becomes
-$E(\mathbf{x})=-\text{log}\sum_{y^{\prime}}\text{exp}\big{(}-E(\mathbf{x},y^{\prime})\big{)}=-\text{log}\sum_{y^{\prime}}\text{exp}\big{(}f(\mathbf{x})[y^{\prime}]\big{)}.$ (3)
+where $\mathbf{h}_{i}^{(0)}=\mathbf{x}_{i}$, and $\mathbf{h}_{i}^{(l+1)}\in\mathbb{R}^{m}$ is the representation of node $v_{i}$ in the $(l+1)$-th layer.
+$\text{Agg}(\cdot)$ computes the incoming message from a neighbor $v_{j}$, and $\text{Upd}(\cdot)$ updates the representation of node $v_{i}$.
+Rather than relying on the raw adjacency matrix $\mathbf{A}$, most GNNs adopt a modified propagation operator $\mathcal{A}$.
+For example, GAT [41] replaces each non-zero entry of $\mathbf{A}$ with a learned attention coefficient that depends on the representations of the corresponding node pair.
+3.3 Node-level Homophily and Heterophily
+For a set of nodes with labels, the homophily ratio of each node quantifies the tendency of the node to share the same label as its neighbors.
+Considering a node $v_{i}$ and its set of neighbors $\mathcal{N}_{i}$, the homophily ratio $h_{i}^{+}$ of $v_{i}$ is defined as: $h_{i}^{+}=\frac{|\{\mathbf{y}_{i}=\mathbf{y}_{j}|v_{j}\in\mathcal{N}_{i}\}|}{|\mathcal{N}_{i}|}$.
+The value of $h_{i}^{+}$ lies in the range $[0,1]$, where values closer to 1 indicate a higher degree of homophily (or lower heterophily), while values nearer to 0 signify the opposite.
+To quantify the homophily of the entire graph $\mathcal{G}$, we compute the average homophily across all nodes: $\mathcal{H}(\mathcal{G})=\frac{\sum_{i=1}^{N}h_{i}^{+}}{N}$.
 4 Methodology
 In this section, we propose a novel graph rewiring strategy that unfolds in three key stages: (i) computing edge torques, (ii) rewiring propagation matrix, and (iii) adjusting message passing.
-The full algorithmic pseudocode is provided in Appendix A, while a detailed complexity analysis of TorqueGNN is included in Appendix B.
+The full algorithmic pseudocode is provided in Appendix B.
 4.1 Derive Graph Torque
@@
 $\mathbf{T}=\mathbf{r}\times\mathbf{F},\quad|\mathbf{T}|=|\mathbf{r}||\mathbf{F}|\sin\theta,$ (2)
-where $\mathbf{r}$ denotes the lever arm vector connecting the point of application to the axis of rotation, $\mathbf{F}$ indicates the force vector, and $\theta$ is the angle between them.
+where $\mathbf{r}$ denotes the displacement vector, $\mathbf{F}$ indicates the force vector, and $\theta$ is the angle between them.
 The magnitude of the torque directly governs an object’s tendency to rotate under the applied force.
-In GNNs, the systematic exploration of node interactions mirrors the lever arm–force relationship underlying torque in classical mechanics.
-Specifically, we can transpose torque to graphs by treating the distance $D_{\langle i,j\rangle}$ between a central node $v_{i}$ and its neighbor $v_{j}$ as a scalar “lever arm”, and the neighbor’s influence level $E_{\langle i,j\rangle}$ as a scalar “disturbing force” on $v_{i}$.
-In graphs, the notion of a disturbing force is abstract and lacks a well-defined direction, as message passing in GNNs does not involve physical "rotation" in the classical mechanical sense.
-Therefore, rather than computing the full vector torque via a cross product, we adopt a scalar approximation based on the product of distance and energy.
-Mathematically, for the edge $k$ connecting nodes $v_{i}$ and $v_{j}$, we define its torque value as follow,
-$T_{e_{k}}=D_{\langle i,j\rangle}\times E_{\langle i,j\rangle},$ (5)
-which measures the disturbance strength that the message carried by edge $e_{k}$ imposes on node $v_{i}$.
-Obviously, larger distances or higher disturbance both amplify torque, while edges maximizing both factors yield the greatest torque that represents the highest priority for graph rewiring.
-Crucially, the definition of $D$ and, in particular, $E$ in Eq.
-(5), remains a key challenge for accurate rewiring.
-Here, we employ an EBM to quantify node interference, and subsequently delineate the construction of both $D$ and $E$.
-Metric 1: Distance.
-To mitigate the effect of noise in raw graphs and features, we estimate node distances using optimized representations.
-Specifically, for a central node $v_{i}$ and its neighbor $v_{j}$, we have the following formulation
-$D_{\langle i,j\rangle}=\|\mathbf{h}_{i}-\mathbf{h}_{j}\|_{2},$ (6)
+In GNNs, the systematic exploration of node interactions can mirror the lever arm–force relationship underlying torque in classical mechanics.
+Specifically, torque on a graph may be conceptualized by treating the displacement vector $\mathbf{D}_{\langle i,j\rangle}$ between a central node $v_{i}$ and its neighbor $v_{j}$ as the effective “lever arm”, while the neighbor’s features $\mathbf{x}_{j}$ act as the applied “force".
+However, the contribution of this force varies across different central nodes.
+Recent studies [42, 7] demonstrate that the generalization of GNNs is influenced by two key factors: the proximity of aggregated features and the disparity in homophily ratio, with smaller values yielding better generalization.
+Inspired by this, we introduce the homophily ratio disparity term $E_{\langle i,j\rangle}$ to modulate the effective force, thereby not only capturing the heterogeneous influences of neighboring nodes but also unifying these two factors within the torque formulation to reduce generalization error.
+Mathematically, for an edge $k$ connecting nodes $v_{i}$ and $v_{j}$, we define the corresponding torque as follow,
+$\mathbf{T}_{e_{k}}=\mathbf{D}_{\langle i,j\rangle}\times E_{\langle i,j\rangle}\mathbf{x}_{j}.$ (3)
+Its magnitude, denoted $T_{e_{k}}$, measures the disturbance imposed by the message passing along edge $e_{k}$ on node $v_{i}$.
+The value naturally increases with larger distances or higher structural disparity, with edges maximizing both factors yielding the greatest torque value that represents the highest priority for graph rewiring.
+A central goal is therefore to provide a principled definition of the displacement vector $\mathbf{D}_{\langle i,j\rangle}$ and the homophily ratio disparity $E_{\langle i,j\rangle}$ in Eq.
+3, after which we detail their construction.
+Metric 1: Displacement Vector.
+To mitigate the effect of noise in raw graphs and features, we estimate the displacement vector $\mathbf{D}_{\langle i,j\rangle}$ using optimized node representations and compute the pairwise distances as
+$\mathbf{D}_{\langle i,j\rangle}=\mathbf{h}_{i}-\mathbf{h}_{j},D_{\langle i,j\rangle}=\|\mathbf{h}_{i}-\mathbf{h}_{j}\|_{2},$ (4)
 […]
-Metric 2: Energy.
-While graph-based message passing can markedly enhance node representations, it also risks amplifying structural and feature noise.
-To mitigate this, we posit nodes that persistently absorb noisy signals exhibit elevated interference strength relative to benign nodes.
-EBM, which assigns scalar energy values to inputs and has demonstrated remarkable robustness against adversarial perturbations [45, 46], provide a foundation for this purpose.
-Therefore, we introduce an EBM-derived energy score as another indicator of edge reliability, as shown in Eq.
-(3), which assigns low energy to benign samples and high energy to corrupted ones.
-In our framework, an energy-based GNN jointly performs node classification and energy estimation.
-For a network parameterized by $\mathbf{\Theta}$, the energy of node $v_{i}$ is defined as $E_{\mathbf{\Theta}}(\mathbf{x}_{i})=-\text{log}\sum_{y}\text{exp}(\mathbf{h}_{i}[y]\big{)}$.
-Here, $\mathbf{h}_{i}$ denotes the logit vector for node $v_{i}$.
-We then derive an edge-level energy metric by combining the energy scores of node pairs,
-$E_{\langle i,j\rangle}=E_{\mathbf{\Theta}}(\mathbf{x}_{i})\times E_{\mathbf{\Theta}}(\mathbf{x}_{j}).$ (7)
+Metric 2: Homophily Ratio Disparity.
+Considering that recent studies emphasize the importance of capturing the homophily ratio disparity in addressing heterophilous graphs, we weight the neighbor features using this disparity to incorporating both it and distance into the torque formulation.
+To estimate node-level homophily, it is crucial to annotate the labels of neighboring nodes around a given node.
+Since labeled data are often scarce, we utilize the model’s final outputs to generate pseudo-labels for unlabeled nodes, with their accuracy improving as the model is progressively optimized.
+Formally, $E_{\langle i,j\rangle}$ is computed by
+$E_{\langle i,j\rangle}=|h_{i}^{+}-h_{j}^{+}|,\ h_{i}^{+}=\frac{|\{v_{j}|\hat{\mathbf{y}}_{i}=\hat{\mathbf{y}}_{j},v_{j}\in\mathcal{N}_{i}\}|}{|\mathcal{N}_{i}|}.$ (5)
+Here, $\hat{\mathbf{y}}_{i}$ denotes the ground-truth label for labeled nodes or the pseudo-label for unlabeled nodes. Finally, the torque value of edge $e_{k}$ is computed by:
+$T_{e_{k}}=\|\mathbf{D}_{\langle i,j\rangle}\times(E_{\langle i,j\rangle}\mathbf{h}_{j})\|_{2}=\sqrt{D_{\langle i,j\rangle}^{2}\cdot(E_{\langle i,j\rangle}\|\mathbf{h}_{j}\|_{2})^{2}-(E_{\langle i,j\rangle}\mathbf{D}_{\langle i,j\rangle}\cdot\mathbf{h}_{j})^{2}}$ (6)
+This formulation captures the combined effects of distance and disparity, facilitating a physics-inspired approach to graph rewiring.
 4.2 Adjust Message Passing
@@
 Herein, we propose an automated threshold learning mechanism that identifies the optimal number of edges to prune by pinpointing the largest successive torque gap.
-Specifically, we first rank all $K$ edges in descending order of their torque values to form a torque-sorted list (TSL), denoting its $k$-th entry entry as $\widetilde{e}_{k}$ with torque $\widetilde{T}_{e_{k}}$, so that $\widetilde{T}_{e_{1}}\geq\widetilde{T}_{e_{2}}\geq\cdots\geq\widetilde{T}_{e_{K}}$.
+Specifically, we first rank all $K$ edges in descending order of their torque values to form a torque‐sorted list (TSL), denoting its $k$-th entry as $\widetilde{e}_{k}$ with torque $\widetilde{T}_{e_{k}}$, so that $\widetilde{T}_{e_{1}}\geq\widetilde{T}_{e_{2}}\geq\cdots\geq\widetilde{T}_{e_{K}}$.
 We then calculate the torque gap between two consecutive links by
 $G_{k,k+1}=\mu_{k}\times\frac{\widetilde{T}_{e_{k}}}{\widetilde{T}_{e_{k+1}}+\delta},$ (7)
-where $\delta$ is a small constant to prevent division by zero, and the weight $\mu_{k}$ reflects the proportion of anomalous edges, those whose distance $D$, energy $E$ and torque $T$ all exceed their respective means, that are captured within the top $k$ torque-ranked set, emphasizing the boundary between desirable and undesirable connections.
+where $\delta$ is a small constant to prevent division by zero, and the weight $\mu_{k}$ reflects the proportion of anomalous edges, those whose distance $D$, disparity $E$ and torque $T$ all exceed their respective means, that are captured within the top $k$ torque-ranked set, emphasizing the boundary between desirable and undesirable connections.
 The computation formula of $\mu_{k}$ is defined as
 $\begin{split}\mu_{k}&=\frac{|High\_e\cap Top\_k|}{|High\_e|},\\ \hfill High\_e&=\{e_{k}\doteq{\langle i,j\rangle}|D_{\langle i,j\rangle}\geq\bar{D},E_{\langle i,j\rangle}\geq\bar{E},T_{\langle i,j\rangle}\geq\bar{T},\langle i,j\rangle\in\mathcal{E}\},\\ Top\_k&=\{e_{k}\doteq{\langle i,j\rangle}|Top_{k}\{T_{\langle i,j\rangle}\},\langle i,j\rangle\in\mathcal{E}\},\end{split}$ (8)
-where $\bar{D},\bar{E},\bar{T}$ denote the mean values of distance, energy and torque, respectively, computed over all $K$ edges.
-The set $High\_e$ comprises edges exhibiting above-average values across all three metrics, while $Top\_k$ contains the top $k$ connections in TSL.
-According to (8), we can identify the optimal cutoff by locating the largest torque gap $\mathcal{K}=\arg\max\limits_{0\leq k\leq K-1}G_{k,k+1}$, which separates the edge set into two groups: undesirable connections ($\widetilde{e}_{1},\cdots,\widetilde{e}_{\mathcal{K}}$) and desirable connections ($\widetilde{e}_{\mathcal{K}+1},\cdots,\widetilde{e}_{K}$).
-In practice, multi-layer GNNs, such as APPNP [47], are widely adopted to enlarge the receptive field of graph convolutions.
+where $\bar{D},\bar{E},\bar{T}$ denote the mean values of distance, disparity and torque, respectively, computed over all $K$ edges.
+The set High_eHigh\_e comprises edges exhibiting above-average values across all three metrics, while Top_kTop\_k contains the top $k$ connections in TSL.
+According to Eq.
+7, we can identify the optimal cutoff by locating the largest torque gap $\mathcal{K}=\arg\max\limits_{0\leq k\leq K-1}G_{k,k+1}$, which separates the edge set into two groups: undesirable connections ($\widetilde{e}_{1},\cdots,\widetilde{e}_{\mathcal{K}}$) and desirable connections ($\widetilde{e}_{\mathcal{K}+1},\cdots,\widetilde{e}_{K}$).
+In practice, multi-layer GNNs, such as APPNP [43] and GCNII [27], are widely adopted to enlarge the receptive field of graph convolutions.
+To allow each layer to adapt the adjacency relationships based on the current information and focus on different structural features of the graph, we adopt a hierarchical rewiring strategy.
 Building on the pairwise torque gap formulation introduced above, we extend this mechanism across multiple propagation layers.
-In specific, for each order $l$, we construct a dedicated propagation matrix that enables selective filtering of undesirable high-order interactions.
+In specific, for each layer $l$, we construct a dedicated propagation matrix that enables selective filtering of undesirable high-order interactions.
 Let $\mathcal{A}^{(l)}$ denote the refined propagation matrix used in $l$-th layer and $\mathcal{A}^{(0)}=\mathbf{A}$.
-We have current neighbors $v_{j}$ of a given node $v_{i}$ based on $\mathcal{A}^{(l)}$, and recompute two key properties as follows
-$D_{\langle i,j\rangle}^{(l+1)}=\|\mathbf{h}_{i}^{(l)}-\mathbf{h}_{j}^{(l)}\|_{2},\quad E_{\langle i,j\rangle}^{(l+1)}=E_{\mathbf{\Theta}}^{(l)}(\mathbf{x}_{i})\times E_{\mathbf{\Theta}}^{(l)}(\mathbf{x}_{j}),$ (10)
-[…]
-[…]
+We identify the current neighbors $v_{j}$ of a given node $v_{i}$ based on $\mathcal{A}^{(l)}$, and recompute the torque as follows
+$\mathbf{T}_{\langle i,j\rangle}^{(l+1)}=(\mathbf{h}_{i}^{(l)}-\mathbf{h}_{j}^{(l)})\times E_{\langle i,j\rangle}\mathbf{h}_{j}^{(l)},$ (9)
+where $l=0,\cdots,L-1$.
+Consequently, we gain the (l+1)(l+1)‑th order torque $T^{(l+1)}$ and the corresponding gap $G^{(l+1)}$ using Eqs.
+6-8, from which we derive a pruned propagation matrix $\mathcal{A}^{(l+1)^{*}}$ with $(K-\mathcal{K})$ non-zero elements.
+Edge-addition High-order Rewiring.
+In the previous steps, we remove undesirable neighbors by computing the torque of existing edges based on two key attributes.
+Extending this strategy, we also consider expanding the receptive field by adding edges that are initially absent but potentially beneficial for message passing.
+However, evaluating torque across all missing edges is computationally intractable, so that we construct a candidate set $\mathcal{T}$ by selecting, for each node, its top-$t$ most similar peers.
+We then compute the $(l+1)$-th order torque $T^{(l+1)}$ for the resulting $N\times t$ candidate edges, and select $r\times N\times t$ edges with the lowest torque values, where $r$ is a sampling ratio.
+Nevertheless, this hard selection process is inherently non-differentiable and thus cannot be used in gradient-based optimization.
+To overcome this, we adopt the Gumbel-Softmax reparameterization trick [44], which enables differentiable sampling by approximating discrete decisions with a continuous relaxation.
+For each candidate edge $k$, we define its logits $\bm{\pi}_{k}=[\pi_{k0},\pi_{k1}]$, where $\pi_{k0}=T_{e_{k}}^{(l+1)}$ (discard) and $\pi_{k1}=1-T_{e_{k}}^{(l+1)}$ (select).
+Drawing independent noise $g_{kj}\sim\text{Gumbel (0, 1)}$, the soft selection probabilities are computed via
 $p_{kj}=\frac{\exp\left(\frac{\log\left(\pi_{kj}\right)+g_{kj}}{\tau}\right)}{\sum_{m=1}^{2}\exp\left(\frac{\log\left(\pi_{km}\right)+g_{km}}{\tau}\right)},\forall j=0,1,k\in\{1,2,\ldots,N\times t\},$ (10)
@@
 Finally, we construct the rewired propagation matrix $\mathcal{A}^{(l+1)}$ by augmenting $\mathcal{A}^{(l+1)^{*}}$ with these probabilistically weighted candidate edges, followed by the standard renormalization procedure.
-[…]
+Messaging Passing on Rewired Graph.
+To avoid misleading representations in the early stages of training, which could either discard important neighbor information or propagate erroneous signals, rewiring at each layer is always performed with respect to the original input graph.
+By rewiring the adjacency matrix $\mathbf{A}$ as described, each propagation layer is endowed with an expanded receptive field capable, enabling the capture of effective multi-level interactions.
+To evaluate the effectiveness of the proposed THR in capturing high-order information in multi-layer GNNs, we use the deep-based, APPNP, as an example.
+Subsequent ablation studies and parameter analyses are conducted within this framework.
+Let $\mathcal{N}_{i}^{(l+1)}=\{v_{j}|\mathcal{A}_{\langle i,j\rangle}^{(l+1)}\neq 0\}$ denotes the refined $(l+1)$-layer neighborhood of node $v_{i}$; then the forward propagation at the $(l+1)$-th layer of APPNP can be reformulated as:
 $\mathbf{h}_{i}^{(l+1)}$ $=\text{ReLU}\Big(\sum\limits_{j\in\{v_{i}\}\cup\mathcal{N}_{i}^{(l+1)}}\alpha\mathbf{h}_{j}^{(l)}+(1-\alpha)\mathbf{h}_{i}^{(0)}\Big).$ (11)
-The final node representations from the last layer are fed into a fully connected layer parameterized by $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ to yield the predicted class probabilities.
-These predictions are compared against the ground-truth labels using a cross-entropy loss minimized through gradient-based optimization.
-4.3 Energy Calibration Objective
-The energy function represents the unnormalized negative log-likelihood of node $v_{i}$, where a lower energy corresponds to a higher likelihood and, consequently, a greater probability of the sample being benign.
-In EBMs, the fundamental objective is to assign low energy to positive samples and high energy to others, thereby maximizing the likelihood of the observed data.
-This is typically achieved by minimizing the negative log-likelihood, formulated as
-$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\text{log}p_{\mathbf{\Theta}}(\mathbf{x}_{i}).$ (13)
-[…]
-$\begin{split}\mathcal{F}\big{(}p_{data}(\mathbf{x})\|p_{\mathbf{\Theta}}(\mathbf{x})\big{)}=\mathbb{E}_{p_{data}(\mathbf{x})}\Big{[}\frac{1}{2}\|\nabla_{\mathbf{\mathbf{x}}}\log p_{data}(\mathbf{x})-\nabla_{\mathbf{x}}\log p_{\mathbf{\Theta}}(\mathbf{x})\|^{2}\Big{]}.\end{split}$ (14)
-To enlarge the energy signal of undesirable connections to facilitate more clearer distinction during edge selection, inspired by [46], we generate negative examples by randomly perturbing the graph structure, i.e., by adding or removing edges.
-Let $\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{i})$ denote the energy of node $v_{i}$ on this perturbed graph.
-We then define an energy-based score according to the following definition
-Definition 1
-(Energy Score): For a node $v_{i}$, its score $S_{\mathbf{\Theta}}(\mathbf{x}_{i})$ assigned by the energy model is surrogated by a gradient in the discrete space:
-$S_{\mathbf{\Theta}}(\mathbf{x}_{i})=\nabla_{\mathbf{x}_{i}}E_{\mathbf{\Theta}}(\mathbf{x}_{i})=\Big{[}\frac{E_{\mathbf{\Theta}}(\mathbf{x}_{1})-\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{1})}{E_{\mathbf{\Theta}}(\mathbf{x}_{1})},\cdots,\frac{E_{\mathbf{\Theta}}(\mathbf{x}_{M})-\widehat{E}_{\mathbf{\Theta}}(\mathbf{x}_{M})}{E_{\mathbf{\Theta}}(\mathbf{x}_{M})}\Big{]},$ (15)
-where $M$ is the number of perturbed samples used. In fact, since $-\log p_{\mathbf{\Theta}}(\mathbf{x})=E_{\mathbf{\Theta}}(\mathbf{x})$, Eq. (15) equals to the following formula,
-$\begin{split}S_{\mathbf{\Theta}}(\mathbf{x}_{i})=\nabla_{\mathbf{x}_{i}}E_{\mathbf{\Theta}}(\mathbf{x}_{i})=\Big{[}\frac{\log p_{\mathbf{\Theta}}(\mathbf{x}_{1})-\log\widehat{p}_{\mathbf{\Theta}}(\mathbf{x}_{1})}{\log p_{\mathbf{\Theta}}(\mathbf{x}_{1})},\cdots,\frac{\log p_{\mathbf{\Theta}}(\mathbf{x}_{M})-\log\widehat{p}_{\mathbf{\Theta}}(\mathbf{x}_{M})}{\log p_{\mathbf{\Theta}}(\mathbf{x}_{M})}\Big{]}.\end{split}$ (16)
-Due to the energy score surrogate, the objective (14) of score matching is reformulated as follows
-$\mathcal{F}\big{(}p_{data}(\mathbf{x}_{i})\|p_{\mathbf{\Theta}}(\mathbf{x}_{i})\big{)}=\mathbb{E}_{p_{data}(\mathbf{x}_{i})}\Big{[}\frac{1}{2}\|S_{data}(\mathbf{x}_{i})-S_{\mathbf{\Theta}}(\mathbf{x}_{i})\|^{2}\Big{]}.$ (17)
-However, the real data is unavailable during training, and thus following [49, 46] and Definition 1, the final optimization formula is given by
-$\mathcal{L}_{sm}=\frac{1}{N}\sum_{i=1}^{N}\Big{[}\nabla_{\mathbf{x}_{i}}S_{\mathbf{\Theta}}(\mathbf{x}_{i})^{\top}\nabla_{\mathbf{x}_{i}}S_{\mathbf{\Theta}}(\mathbf{x}_{i})+\frac{1}{2}\|S_{\mathbf{\Theta}}(\mathbf{x}_{i})\|^{2}\Big{]}.$ (18)
-By minimizing the surrogate loss (18), we not only approximate the original score matching objective of the EBM, but also explicitly amplify the energy signals associated with undesirable edges.
+Here, $\alpha$ controls the trade-off between the hidden representation and the residual connection.
+The initial representation $\mathbf{h}_{i}^{(0)}=\mathbf{x}_{i}\mathbf{\Theta}$ is computed through a linear transformation of the input feature $\mathbf{x}_{i}$.
+The final node representations from the last layer are passed through a fully connected layer parameterized by $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$, which yields the predicted class probabilities.
+These predictions are compared against the ground-truth labels using a cross-entropy loss, which is minimized through gradient-based optimization.
+5 Complexity Analysis
+The dominant computational cost of THR lies in: 1) Torque computation and graph rewiring.
+For each order $l$, we compute torque values only on the edges in $\mathcal{A}^{(l)}$, costing $\mathcal{O}(|\mathcal{A}^{(l)}|)$, and then sort these values in $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$.
+When adding edges, if the candidate set size is $B$, the combined probability calculation and sorting cost is $\mathcal{O}(B+B\log B)$.
+2) Message passing on the rewired graph $\mathcal{A}^{(l)}$.
+For the input layer with parameter $\mathbf{\Theta}\in\mathbb{R}^{d\times m}$ on $\mathbf{X}\in\mathbb{R}^{N\times d}$, it costs $\mathcal{O}(Ndm)$.
+Aggregation over $\mathcal{A}^{(l)}$ then costs $\mathcal{O}(m|\mathcal{A}^{(l)}|)$ per layer.
+The output layer with $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ requires $\mathcal{O}(Nmc)$.
+Putting these together for an $L$-layer network and assuming $B\ll|\mathcal{A}^{(l)}|$ for all $l$, the overall complexity is $\mathcal{O}(Ndm+\sum_{l=1}^{L}|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$, which is slightly higher than that of standard methods with $\mathcal{O}(Ndm+m|\mathcal{A}^{(l)}|)$.
 6 Experiments
 Datasets.
-We evaluate our method on ten standard node classification benchmarks, including three homophilous graphs (Citeseer, Cora, and Pubmed) and seven heterophilous graphs (Texas, Wisconsin, Cornell, Actor, Tolokers, Questions, and Penn94).
-Among them, Tolokers, Questions, and Penn94 are large-scale datasets.
-Following previous work, we adopt the same data split strategy: 48% of the nodes are used for training, 32% for validation, and the remaining 20% for testing.
-The statistics of these datasets are summarized in Table 1, and further details can be found in Appendix C.2.
+We evaluate our method on eleven standard node classification benchmarks, which include six heterophilic datasets: Texas, Wisconsin, Cornell, Actor, Penn94 and Flickr; five homophilous graphs Citeseer, Cora, Pubmed, Tolokers and Questions.
+Among them, Tolokers, Questions, Penn94 and Flickr are large-scale datasets.
+The statistics for these datasets are summarized in Table 1, with further details provided in Appendix C.2.
 Table 1: Benchmark dataset statistics.
@@
 Actor 0.22 7,600 26,752 5 931
+Citeseer 0.74 3,327 4,676 7 3,703
+Cora 0.81 2,708 5,278 6 1,433
+Pubmed 0.80 19,717 44,327 3 500
 Tolokers 0.60 11,758 51,900 2 10
@@
 Penn94 0.47 41,554 1,362,229 2 4,814
-Citeseer 0.74 3,327 4,676 7 3,703
-Cora 0.81 2,708 5,278 6 1,433
-Pubmed 0.80 19,717 44,327 3 500
-Competitors.
-To evaluate the effectiveness of our model in capturing high-order interactions, we adopt APPNP as the downstream GNN, due to its suitability for deep graph architectures.
-We compare our approach against eleven baselines: a two-layer MLP; three classic homophily-oriented GNNs, including GCN [50], GAT [41] and APPNP [47]; and seven representative methods specifically designed for heterophilous graphs: H2GCN [51], GPR-GNN [52], FAGCN [39], ACM-GNN [38], OGNN [53], M2M-GNN [40], PCNet [54].
-Furthermore, to assess the impact of our THR rewiring strategy, we compare it with other rewiring strategies: first-order spectral rewiring (FoSR) [55], joint denoising and rewiring (JDR) [36] and randomly removing edges (DropEdge) [56].
-More details on all methods are provided in Appendix C.3.
+Flickr 0.32 89,250 2,724,458 7 500
+Baselines.
+THR is a plug-in module that can be integrated into various state-of-the-art GNNs.
+To evaluate the improvements offered by THR for GNNs, we select three representative models for experimentation, including two models designed for homophilous graphs: the vanilla GCN [45] and the deep-based APPNP [43], and GPRGNN [46] that is designed for heterophilous graphs.
+To evaluate the effectiveness of THR in comparison to other graph rewiring techniques, we select five superior methods, including: First-order Dpectral Rewiring (FoSR) [32], Batch Ollivier-Ricci Flow (BORF) [47], Stochastic Jost and Liu Curvature Rewiring (SJLR) [48], Deep Heterophily Graph Rewiring (DHGR) [14] and randomly edge removal (DropEdge).
+Here, we adopt layer-wise DropEdge (Dropedge-L), as proposed by [49], to ensure a fair comparison with the hierarchical structure of THR.
+Further details on all methods are provided in Appendix C.3.
 Setups.
-We report node classification ACC, defined as the proportion of correctly predicted labels.
-For most benchmark datasets, models are trained using the Adam optimizer for 100 epochs with a learning rate of 0.001 and 512 hidden units.
-For larger datasets such as Tolokers, Questions, and Penn94, the number of training epochs is increased to 500 to ensure sufficient training.
-Detailed hyperparameter and environment configurations are provided in Appendix C.4, including the code link in C.1.
-Each experiment is repeated ten times, and we report the mean and standard deviation.
+We report node classification accuracy (ACC), defined as the proportion of correctly predicted labels.
+For all benchmark datasets, models are trained using the Adam optimizer.
+Competitors are performed based on their respective source code.
+Detailed hyperparameters and environment configurations for THR are provided in Appendix C.4.
+Following prior work [37, 50], we adopt the data split strategy for all methods: 48% of the nodes are used for training, 32% for validation, and the remaining 20% for testing.
+Each experiment is conducted over 10 runs with different random splits, and the results are reported as the mean and standard deviation.
+Table 2: Node classification results on benchmark datasets with GCN and GPRGNN as the backbone models: Mean ACC % (Standard Deviation %).
+The first- and second-best accuracies are highlighted in red and green, respectively.
+Methods/Datasets Citeseer Cora Pubmed Texas Wisconsin Actor Cornell
+GCN 75.52 (2.19) 86.96 (1.27) 86.43 (0.38) 58.61 (7.18) 52.60 (8.72) 30.15 (1.03) 57.50 (4.66)
+FoSR 78.03 (1.45) 87.00 (1.21) 86.34 (0.31) 74.70 (6.23) 65.58 (4.89) 30.16 (1.03) 54.59 (5.01)
+BROF 78.45 (1.52) 86.86 (1.35) 86.42 (0.38) 74.51 (6.26) 65.59 (4.52) 30.20 (1.17) 60.27 (3.64)
+SJLR 77.87 (1.81) 86.60 (1.64) 86.52 (1.73) 60.14 (0.89) 55.16 (0.95) 30.80 (1.34) 58.11 (6.86)
+DHGR 78.68 (1.51) 86.61 (1.73) 86.40 (0.38) 60.20 (6.39) 66.07 (12.51) 34.39 (0.99) 58.68 (5.01)
+DropEdge-L 74.93 (1.85) 86.62 (1.23) 83.07 (2.58) 62.74 (8.32) 58.82 (8.24) 32.97 (0.92) 54.32 (3.72)
+THR 80.43 (1.52) 86.97 (1.19) 87.21 (0.45) 76.27 (4.67) 68.09 (2.71) 33.20 (0.90) 58.91 (9.11)
+GPRGNN 77.37 (1.83) 87.34 (1.14) 87.21 (0.43) 89.22 (5.56) 87.94 (5.29) 37.27 (1.16) 80.27 (6.63)
+FoSR 77.37 (1.83) 87.52 (1.63) 87.22 (0.46) 90.20 (5.04) 89.85 (3.45) 37.25 (1.19) 84.05 (7.88)
+BORF 78.77 (1.67) 87.49 (1.24) 87.17 (0.39) 91.16 (5.15) 89.11 (4.32) 37.52 (1.06) 85.49 (4.83)
+SJLR 78.38 (1.49) 86.97 (1.63) 88.11 (0.41) 90.00 (2.83) 89.26 (6.38) 34.87 (1.69) 81.62 (9.35)
+DHGR 77.77 (2.06) 87.19 (1.39) 87.69 (0.47) 89.02 (4.31) 86.03 (6.32) 35.20 (1.20) 84.31 (4.56)
+DropEdge-L 78.73 (1.91) 86.91 (1.07) 87.50 (0.48) 90.17 (3.06) 87.79 (6.28) 37.77 (1.16) 84.05 (9.00)
+THR 79.15 (1.69) 87.60 (1.15) 88.28 (0.52) 91.96 (3.76) 91.91 (4.75) 38.00 (0.56) 86.22 (5.19)
 Node Classification Results.
-Table 2 reports test-set accuracy for three TorqueGNN variants, including edge-addition THR (A-THR), edge-removal THR (R-THR) and mixed THR (M-THR), alongside eleven state-of-the-art baselines across seven benchmark datasets.
-Several insights emerge: 1) On every dataset, at least one TorqueGNN variant appears among the top three performers, and TorqueGNN achieves the highest accuracy on the majority of benchmarks.
-2) Three rewiring strategies yield substantial gains over APPNP and improvements are most pronounced on heterophilous graphs, underscoring the significance of our torque-based rewiring.
-3) On Citeseer, Cora, and Pubmed, both A-THR and R-THR outperform the M-THR variant, suggesting that simultaneous addition and removal can inadvertently constrain information flow for citation networks.
-4) In highly heterophilous datasets, although edge addition boosts APPNP, original undesirable links limit its effect.
-By contrast, R-THR achieves more high accuracy by judiciously deleting these connections.
-5) In cases such as Pubmed and Cornell, R-THR attains the best results, likely because aggressive removal of noisy original edges without introducing additional candidates yields a cleaner propagation structure.
-Table 2: Node classification results on benchmark datasets: Mean ACC % ± Standard Deviation %. he first-, second- and third-best accuracy are highlighted in red, green and blue, respectively.
-Methods Citeseer Cora Pubmed Texas Wiscon. Cornell Actor
-MLP 74.02±1.90 75.69±2.00 87.16±0.37 80.81±4.75 85.29±3.31 81.89±6.40 36.53±0.70
-GCN 76.50±1.36 86.98±1.27 88.42±0.50 55.14±5.16 51.76±3.06 60.54±5.30 27.32±1.10
-GAT 76.55±1.23 87.30±1.10 86.33±0.48 52.16±6.63 49.41±4.09 61.89±5.05 27.44±0.89
-H2GCN 77.11±1.57 87.87±1.20 89.49±0.38 84.86±7.23 87.65±4.98 82.70±5.28 35.70±1.00
-GPR-GNN 77.13±1.67 87.95±1.18 87.54±0.38 78.38±4.36 82.94±4.21 80.27±8.11 34.63±1.22
-FAGCN 74.01±1.85 86.34±0.67 76.57±1.88 77.56±6.11 79.41±6.55 78.64±5.47 34.85±1.61
-ACM-GCN 77.32±1.70 87.91±0.95 90.00±0.52 87.84±4.40 88.43±3.22 85.14±6.07 36.28±1.09
-OGNN 77.31±1.73 88.37±0.75 90.15±0.38 86.22±4.12 88.04±3.63 87.03±4.73 37.99±1.00
-M2M-GNN 77.20±1.80 88.12±1.00 90.35±0.60 89.19±4.50 89.01±4.10 86.48±6.10 36.72±1.60
-PCNet 77.50±1.06 88.41±0.66 89.51±0.28 88.11±2.17 88.63±2.75 82.61±2.70 37.80±0.64
-APPNP (backbone) 74.02±0.38 88.32±0.41 84.53±0.34 49.17±3.30 47.60±4.54 55.83±4.02 29.03±1.12
-TorqueGNN w M-THR 75.86±0.23 84.78±0.28 87.52±0.06 91.11±1.67 89.60±1.20 89.44±1.67 36.72±0.16
-TorqueGNN w R-THR 76.51±0.09 86.73±0.50 90.18±0.24 88.61±0.83 86.20±0.60 90.83±2.50 36.07±0.36
-TorqueGNN w A-THR 78.03±0.05 88.47±0.28 89.38±0.02 86.11±0.01 82.00±0.01 83.61±3.15 36.24±0.04
-Table 3: Node classification results on large-scale datasets: Mean ACC % (ROC AUC for imbalanced Questions) ± Standard Deviation %, where the optimal and suboptimal results are highlighted in red and green, respectively.
-Penn94 Tolokers Questions
-baseGNN 75.19 (0.17) 76.47 (1.16) 46.57 (0.71)
-FoSR 75.21 (0.19) 76.64 (1.15) 46.95 (0.83)
-JDR 75.19 (0.23) 78.06 (0.88) 46.98 (0.87)
-DropEdge 78.46 (0.02) 77.03 (0.26) 60.01 (0.24)
-THR 80.88 (0.01) 78.08 (0.16) 62.64 (0.89)
+Table 2 presents the test-set accuracy gains achieved by various rewiring approaches on GCN and GPRGNN across seven benchmark datasets.
+Several key insights can be drawn: 1) Compared to the baselines, all rewiring methods show performance improvements on most datasets, with particularly notable gains on heterophilous graphs.
+2) In all datasets, the proposed THR ranks among the top two performers, achieving the highest accuracy gain on the majority of benchmarks.
+3) Although FoSR, BORF, and DHGR also exhibit strong performance on certain datasets, their gains are only marginally higher than those of THR.
+Overall, THR outperforms these methods and delivers the best results in all cases when GPRGNN is used as the downstream model.
+4) DropEdge-L, which is also based on hierarchical graph rewiring, outperforms other rewiring methods on some datasets (e.g., Texas and Actor), validating the effectiveness of the hierarchical strategy.
+Although DropEdge shows performance improvements on certain datasets, its inherent randomness negatively impacts the model’s performance, resulting in lower performance than the baseline in some cases, e.g., Citeseer.
+This further validates the effectiveness of the proposed torque-driven hierarchical approach.
+Table 3: Node classification results on large-scale datasets: Mean ACC % (ROC AUC for imbalanced Questions and Tolokers) (Standard Deviation %), where the optimal and suboptimal results are highlighted in red and green, respectively.
+OoM means that the model suffers from the out-of-memory error.
+Methods/Datasets Questions Tolokers Penn94 Flickr
+GCN 75.26 (0.84) 83.79 (0.74) 80.18 (0.36) 57.48 (7.85)
+FoSR 75.19 (0.71) 84.14 (0.99) 80.19 (0.35) 58.03 (6.75)
+BROF 75.15 (0.84) MemoryError OoM OoM
+SJLR 72.07 (6.12) 76.14 (1.14) 80.20 (0.28) 64.49 (2.82)
+DHGR OoM 76.45 (12.16) OoM OoM
+DropEdge-L 74.06 (1.11) 84.00 (0.65) 62.27 (0.35) 59.29 (2.26)
+THR 75.92 (1.09) 84.43 (0.88) 80.32 (0.23) 68.29 (0.78)
+GPRGNN 72.89 (1.42) 71.99 (0.93) 84.18 (0.30) 48.90 (5.81)
+FoSR 72.91 (1.43) 71.99 (0.93) 84.22 (0.29) 49. 16 (6.49)
+BORF 72.99 (1.44) MemoryError OoM OoM
+SJLR 72.27 (1.24) 69.46 (1.07) 83.89 (0.20) 61.61 (3.22)
+DHGR OoM 70.96 (1.14) OoM OoM
+DropEdge-L 72.07 (1.35) 71.98 (1.09) 83.67 (0.44) 63.42 (3.26)
+THR 73.41 (0.98) 72.05 (1.24) 84.45 (0.29) 65.29 (3.30)
 Results on Larger Graphs.
 Scalability of rewiring techniques on large graphs is crucial, particularly for end-to-end methods that dynamically add and remove edges during training.
-In TorqueGNN, the principal computational burden stems from computing high-order torques and corresponding gaps, which incurs a complexity of $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$ (see Appendix B for details).
-Despite this overhead, THR remains tractable for large graphs.
-Table 3 compares multiple rewiring schemes on larger heterophilous datasets: Both THR and DropEdge yield substantial accuracy improvements over the base GNN, yet THR consistently outperforms all alternatives.
-In contrast, FoSR and JDR achieve only marginal gains.
-Moreover, THR surpasses DropEdge, validating the effectiveness of the proposed torque-driven manner.
-Figure 2: Results of TorqueGNN w/o THR and w THR on graphs with noisy edges.
-Results on Adversarially Perturbed Graphs.
-Our torque-driven rewiring not only improves predictive accuracy but also enhances robustness to structural attacks by detecting and suppressing adversarial edges.
-To evaluate this, we employ MetaAttack [57] to inject adversarial links at a perturbation rate of 25% of the original edge count.
-Figure 2 compares classification accuracy under three settings: TorqueGNN without THR (w/o THR) on attacked graph, TorqueGNN with THR (w THR) on attacked graph and TorqueGNN w THR on normal graph.
-In all cases, we employ the mixed addition and deletion strategy for optimization.
-As shown, across both homophilous and heterophilous datasets, applying THR to the perturbed graph substantially improves classification performance, which demonstrates the effectiveness of our method in detecting and eliminating adversarial edges.
-Moreover, on heterophilous datasets, TorqueGNN w THR applied to the attacked graph even surpasses its performance on the normal graph, suggesting that the injected adversarial edges accentuate the distinction between desirable and undesirable connections, thereby enhancing the effectiveness of our rewiring mechanism.
-Figure 3: Ablation study: Performance comparison of GCN and GAT with different rewiring schemes on four datasets.
+In THR, the primary computational cost arises from computing torques and the corresponding gaps, which incurs a complexity of $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$ (see Section 5 for details).
+Despite this overhead, THR remains computationally feasible for large graphs.
+Table 3 compares several rewiring schemes on larger datasets, with THR consistently outperforming all alternatives.
+Notably, with the exception of the Flickr dataset, all rewiring methods show only marginal improvements, and in some cases, even lead to a decline in performance.
+This may be attributed to the fact that the raw graphs of these datasets already contain sufficient structural information, and the rewiring methods introduce only minor modifications, or in some cases, may even result in the loss of critical semantics, thus negatively impacting classification performance.
+Figure 2: Ablation study: Performance comparison of GCN, GPRGNN, and APPNP with various THR variants across six datasets.
 Ablation Study.
-Table 2 has shown that all three THR variants markedly improve upon the APPNP backbone.
-To evaluate their broader applicability, we incorporate each strategy into two classic two-layer GNNs (GCN and GAT), and repeat the ablation studies in Figure 3.
-As illustrated, on the highly heterophilous Texas, Wisconsin, and Actor datasets, every rewiring variant substantially outperforms its base GNN.
-Even on the homophilous Pubmed, R-THR surpasses the base models, echoing the results in Table 2 and underscoring the broad applicability of our torque-driven optimization.
-We note, however, that A-THR and M-THR occasionally underperform relative to the baselines, likely because the current structure is already well-formed, and excessive edge additions may introduce noise, thereby degrading classification performance.
-Ablation study of significant components is included in Appendix C.5.
+We conduct an ablation study to assess the impact of edge removal and addition operations in THR, using GCN, GPRGNN, and APPNP as backbone models.
+THR has three variants: edge-addition THR (A-THR), edge-removal THR (R-THR), and mixed THR (M-THR).
+As illustrated in Figure 2, most rewiring variants significantly outperform their base GNNs.
+However, on the Wisconsin dataset, GPRGNN slightly surpasses GPRGNN w R-THR, likely because GPRGNN effectively allocates sign edges to distinguish class information, while R-THR removes heterophilic links, inadvertently causing the model to lose some discriminative features.
+On the Flickr dataset, R-THR improves performance for all models, while A-THR and M-THR degrade the performance of APPNP.
+Similarly, on PubMed, both A-THR and M-THR reduce the performance of GCN.
+These results suggest that for these graphs, excessive edge addition leads to information interference and confusion of node features, while GPRGNN mitigates this effect by utilizing a sign edge strategy.
+In conclusion, the THR strategy enhances model performance, but its effectiveness varies across datasets with different characteristics.
+[…]
+Table 4: Ablation study: A comparison of THR and its variants by removing specific components. The optimal and suboptimal results are highlighted in bold and underlined, respectively.
+Datasets Texas Wisconsin Cornell Actor Penn94 Flickr
+[…]
+[…]
+[…]
+THR 72.01 (6.13) 75.89 (3.46) 77.30 (7.17) 36.28 (1.15) 76.21 (0.47) 63.28 (1.56)
 Parameter Analysis.
-To assess the influence of the key hyperparameter $\alpha$, we perform a sensitivity analysis shown in Figure 4.
-The figure examine the trade-off parameter $\alpha$, which balances the contribution of the learned hidden representation and the original input features, where a larger $\alpha$ increases the influence of the hidden representations.
-We observe that smaller graphs (e.g., Texas and Cornell), optimal accuracy is achieved at low $\alpha$ value ($0.01\leq\alpha\leq 0.05$), implying that raw node features provide sufficient discriminative power.
-In contrast, on medium-to-large graphs, better performance is observed when $\alpha\geq 0.5$, reflecting the necessity of high-order hidden representations to capture more complex community structures.
+Since the edge-removal procedure automatically determines the cutoff $\mathcal{K}$, we investigate the main hyperparameter $t$ of THR, which defines the number of candidate edges for addition.
+As shown in Figure 3, we present the performance curves for varying $t$ values in $\{2,4,6,8,10\}$ across five datasets.
+On both homophilous and heterophilous datasets, accuracy increases as $t$ grows, demonstrating that the proposed edge-addition scheme aids the model in capturing global information.
+However, this does not imply that adding more edges is always beneficial.
+For instance, on the Flickr dataset, performance decreases when $t=8$, as excessive edge addition may introduce noise, as highlighted in the ablation study.
 Sensitivity analysis of other hyperparameters is presented in Appendix C.5.
-[…]
+Figure 3: Parameter sensitivity: Performance curves on five datasets as the number of candidate edges $t$ varies from 2 to 10.
 7 Conclusion
-In summary, we proposed a Torque-driven hierarchical rewiring graph neural network (TorqueGNN), which dynamically refined the graph structures to enhance representation learning on heterophilous and noisy graphs.
-By introducing an interference-aware torque metric, the product of the distance and the energy between node representations, TorqueGNN automatically removed undesirable connections and introduced beneficial ones during message passing.
-This hierarchical rewiring yielded interference-resilient, importance-aware propagation tailored to each layer’s receptive field.
-Extensive evaluations across homophilous and heterophilous benchmark datasets demonstrated that TorqueGNN consistently outperformed state-of-the-art baselines and maintained robust performance on adversarial graphs.
-We state the limitations of our model in Appendix D.
-Appendix A Algorithm
-Algorithm 1 outlines the complete workflow of TorqueGNN.
-Input: Node features $\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$, candidate edge set $\mathcal{T}$, ground truth matrix $\mathbf{Y}$, hyperparameters $L$ and $\alpha$.
+In summary, we proposed a Torque‑driven hierarchical rewiring strategy (THR), which dynamically refined the graph structures to enhance representation learning on heterophilous and homophilous graphs.
+By introducing an interference‑aware torque metric, the product of the displacement vector and the feature vector weighted by the homophily ratio disparity, THR automatically removed undesirable connections and introduced beneficial ones during message passing.
+This hierarchical rewiring yielded interference‑resilient, importance‑aware propagation tailored to each layer’s receptive field.
+Extensive evaluations across homophilous and heterophilous benchmark datasets demonstrated that THR consistently obtained the performance gains and outperformed other rewiring methods.
+Appendix A Appendix
+Appendix B Algorithm
+Algorithm 1 outlines the complete workflow of APPNP with THR.
+Input: Node features $\{\mathbf{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{N}$, candidate edge set $\mathcal{T}$, ground truth matrix $\mathbf{Y}$, the number of layers $L$, hyperparameters $t$ and $\alpha$.
 Output: The predicted class label.
@@
 […]
-5   Compute pairwise distance $D_{\langle i,j\rangle}^{(l)}$ and energy $E_{\langle i,j\rangle}^{(l)}$ with Eqs. (6) and (7);
-6   Compute the $l$th order torques with Eq. (5) and sort them. // Torque computation
-7   Gain the largest torque gap $\mathcal{K}$ with Eqs. (8)-(9);
+5   Compute pairwise distance $D_{\langle i,j\rangle}^{(l)}$ and homophily disparity $E_{\langle i,j\rangle}$ with Eqs. 4 and 5;
+6   Compute the $l$th order torques with Eq. 6 and sort them. // Torque computation
+7   Gain the largest torque gap $\mathcal{K}$ with Eqs. 7-8;
 8   Remove the top $\mathcal{K}$ edges to gain $\mathcal{A}^{(l)^{*}}$. // Removing undesirable edges
-9   Compute the sampling probability of candidate edges with Eq. (11);
+9   Compute the sampling probability of candidate edges with Eq. 10;
 10   Add beneficial candidate edges to form the refined propagation matrix $\mathcal{A}^{(l)}$. // Adding desirable connections
-11   Update node representation $\mathbf{h}_{i}^{(l)}$ with Eq. (12). // Message passing
+11   Update node representation $\mathbf{h}_{i}^{(l)}$ with Eq. 11. // Message passing
 […]
 […]
-14 for $t=1\rightarrow T$ do
-[…]
-16
-17Obtain $\mathbf{\hat{y}}_{i}=\text{Softmax}\left(\mathbf{h}^{(L)}_{i}\mathbf{\Phi}\right)$;
+14
+15Obtain $\mathbf{\hat{y}}_{i}=\text{Softmax}\left(\mathbf{h}^{(L)}_{i}\mathbf{\Phi}\right)$;
 return The predicted class label of the $i$-th node is given by $\arg\max{\hat{\mathbf{y}}_{i}}$.
-Algorithm 1 TorqueGNN
-Appendix B Complexity Analysis
-The dominant computational cost of TorqueGNN lies in: 1) Torque computation and graph rewiring.
-For each order $l$, we compute torque values only on the edges in $\mathcal{A}^{(l)}$, costing $\mathcal{O}(|\mathcal{A}^{(l)}|)$, and then sort these values in $\mathcal{O}(|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$.
-When adding edges, if the candidate set size is $B$, the combined probability calculation and sorting cost is $\mathcal{O}(B+B\log B)$.
-2) Message passing on the rewired graph $\mathcal{A}^{(l)}$.
-Initial representations $\mathbf{H}^{(0)}$ are obtained by a fully connected layer with parameter $\mathbf{\Theta}\in\mathbb{R}^{d\times m}$ on $\mathbf{X}\in\mathbb{R}^{N\times d}$, at cost $\mathcal{O}(Ndm)$.
-Aggregation over $\mathcal{A}^{(l)}$ then costs $\mathcal{O}(m|\mathcal{A}^{(l)}|)$ per layer.
-The output layer with $\mathbf{\Phi}\in\mathbb{R}^{m\times c}$ requires $\mathcal{O}(Nmc)$.
-Putting these together for an $L$-layer network and assuming $B\ll|\mathbf{A}^{l}|$ for all $l$, the overall complexity is $\mathcal{O}(LNdm+\sum_{l=1}^{L}|\mathcal{A}^{(l)}|\log|\mathcal{A}^{(l)}|)$.
+Algorithm 1 GNN with THR
 Appendix C More Experimental Results
@@
   • Homophilous Datasets. Citeseer, Cora and Pubmed are three citation networks, and they are published in [51]. Specifically,
-    • Citeseer comprises 3,327 publications classified into six categories, with each paper encoded by a 3,703-dimensional binary word-presence vector.
+    • Citeseer comprises 3,327 publications classified into six categories, with each paper encoded by a 3,703-dimensional binary word‐presence vector.
     • Cora consists of 2,708 scientific publications classified into seven research topics.
@@
 Papers are described by 500-dimensional term frequency–inverse document frequency feature vectors, and citation edges capture scholarly references.
+    • Tolokers [52] is built from the Toloka crowdsourcing platform, comprising 11,758 nodes and 519,000 edges that link workers who collaborated on the same task.
+Each node carries a 10-dimensional feature vector and is assigned one of two labels based on whether the worker was banned.
+    • Questions [52] is an interaction graph of users on the Yandex Q question-answering platform, comprising 48,921 nodes and 153,540 edges that link users who interacted on the same question.
+Each node carries a 301-dimensional feature vector and a binary label for node classification.
   • Heterophilous Datasets
@@
     • Actor [53] is the actor-only induced subgraph of a film–director–actor–writer network on Wikipedia, where each node represents an actor and an undirected edge connects two actors if they co-occur on the same Wikipedia page.
-    • Tolokers [61] is built from the Toloka crowdsourcing platform, comprising 11,758 nodes and 519,000 edges that link workers who collaborated on the same task.
-Each node carries a 10-dimensional feature vector and is assigned one of two labels based on whether the worker was banned.
-    • Questions [61] is an interaction graph of users on the Yandex Q question-answering platform, comprising 48,921 nodes and 153,540 edges that link users who interacted on the same question.
-Each node carries a 301-dimensional feature vector and a binary label for node classification.
     • Penn94 [54] is a subgraph of the Facebook100 dataset featuring 41,554 university students as nodes, connected by 1,362,229 undirected friendship edges.
 Each node is described by a five-dimensional feature vector and labeled by the gender of the students.
-C.3 Methods
+    • Flickr [55] is an undirected graph originated from NUS-wide, including 89,250 nodes and 2,724,458 edges.
+Each node is an image with 500-dimensional bag-of-word features and each edge links two images sharing some common properties.
+C.3 Baselines
 C.3.1 GNNs for Homophilous and Heterophilous Graphs
@@
 where $\mathbf{W}$ is the learnable parameter matrix.
-GAT leverages attention mechanisms to weight the influence of neighboring nodes during message passing. Its core rule is:
-$\mathbf{h}_{i}^{(l+1)}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}^{(l)}\right),\alpha_{ij}=\frac{\exp\left(\operatorname{LeakyReLU}\left(a^{\top}\left[\mathbf{W}\mathbf{h}_{i}^{(l)}\|\mathbf{W}\mathbf{h}_{j}^{(l)}\right]\right)\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\operatorname{LeakyReLU}\left(a^{\top}\left[\mathbf{W}\mathbf{h}_{i}^{(l)}\|\mathbf{W}\mathbf{h}_{k}^{(l)}\right]\right)\right)}.$ (20)
 APPNP first achieves the feature transformation by:
@@
 Here, $\mathbf{P}=\mathbf{D}^{-1/2}\mathbf{A}\mathbf{D}^{-1/2}$ is the symmetrically normalized adjacency matrix and $\alpha$ is a trade-off hyperparameter.
-H2GCN contains three designs, and the first design aims to separate ego- and neighbor-message:
-$\mathbf{h}_{i}^{(l)}=[\mathbf{h}_{i}^{(l-1)}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}_{i}\})]$ (23)
-Second, H2GCN directly aggregates neighbors of different hops, which enhances high-order information:
-$\mathbf{h}_{i}^{(l)}=[\mathbf{h}_{i}^{(l-1)}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}^{1}_{i}||\text{Agg}(\{\mathbf{h}_{j}^{(l-1)}:v_{j}\in\mathcal{N}^{2}_{i}||\cdots\})].$ (24)
-Finally, the output is gained by combining the intermediate representations of all layers:
-$\mathbf{h}_{i}=[\mathbf{h}_{i}^{(1)}||\mathbf{h}_{i}^{(2)}||\cdots||\mathbf{h}_{i}^{(L)}].$ (25)
 GPR-GNN generalizes personalized PageRank by treating each hop’s contribution as a learnable parameter:
@@
 where $\gamma^{l}\mathbf{P}$ measures the propagation coefficient for the connection between nodes $v_{i}$ and $v_{j}$.
-FAGCN adaptively balances low- and high-frequency graph signals via a learnable mixing coefficient:
-$\mathbf{h}_{i}^{(l)}=\eta\mathbf{h}_{i}^{(0)}+\sum_{j\in\mathcal{N}_{i}}\frac{\alpha_{ij}}{\sqrt{d_{i}d_{j}}}\eta\mathbf{h}_{i}^{(l-1)},$ (27)
-where $\alpha_{ij}\in(-1,1)$ is a learnable coefficient.
-ACM-GNN dynamically learns how to weight multiple propagation channels:
-$\mathbf{H}^{(l)}=\operatorname{ReLU}\left(\operatorname{diag}\left(\alpha_{H}^{(l)}\right)(\mathbf{I}-\mathbf{P})\mathbf{H}^{(l-1)}\mathbf{W}_{H}^{(l)}+\operatorname{diag}\left(\alpha_{L}^{(l)}\right)\mathbf{P}\mathbf{H}^{(l-1)}\mathbf{W}_{L}^{(l)}+\operatorname{diag}\left(\alpha_{I}^{(l)}\right)\mathbf{H}^{(l-1)}\mathbf{W}_{I}^{(l)}\right).$ (28)
-Here, $\alpha$ is adaptive mixing weight.
-OGNN introduces an order-aware message passing by sorting neighbors according to learned importance scores before aggregation:
-$\mathbf{h}_{i}=\mathbf{h}_{i}^{(1)}\odot g_{i}^{(1)}+\mathbf{h}_{i}^{(2)}\odot\left(g_{i}^{(2)}-g_{i}^{(1)}\right)\ldots+\mathbf{h}_{i}^{(L)}\odot\left(g_{i}^{(KL}-g_{i}^{(L-1)}\right),$ (29)
-where $g_{i}^{(l)}$ is a vector with a split point $p_{i}^{(l)}$ so that its first entries are 0 and its remaining entries are 1.
-M2MGNN partitions neighbor embeddings into class-aligned “chunks” via learnable soft labels:
-$\mathbf{h}_{i}^{(l)}=\operatorname{ReLU}\left((1-\beta)\mathbf{h}_{i}^{(0)}+\beta\mathbf{m}_{i}^{(l)}\right),\mathbf{C}_{it}^{(l)}=\sum_{j\in\mathcal{N}_{i}}\mathbf{s}_{t}^{(l)}(i,j)\hat{\mathbf{h}}_{j}^{(k)},\quad\mathbf{m}_{i}^{(l)}=\|_{t=1}^{C}\mathbf{C}_{it}^{(l)},$ (30)
-where $\mathbf{s}_{t}^{(l)}(i,j)$ is the score of node $v_{i}$ towards $v_{j}$ in the $t$-th chunk.
-PCNet introduces a unified two-fold filtering framework that simultaneously performs homophilous low-pass smoothing and heterophilous high-pass aggregation:
-$\mathbf{H}=(\sum_{k=1}^{K}\theta_{k}\mathrm{PC}_{k}(\tilde{L})+\theta_{\mathrm{0}})\mathbf{X}\mathbf{W}.$ (31)
-Here, $\mathrm{PC}_{k}(\tilde{L})$ denotes the $n$-th Poisson–Charlier polynomial of the normalized Laplacian $\tilde{L}$, and $\{\theta_{k}\}_{k=0}^{K}$ are learnable filter coefficients.
 C.3.2 Rewiring Strategies
-FoSR develope a spectral-expansion–based rewiring algorithm that added edges to alleviate over-squashing, and integrate it with a relational GNN to provably avoid oversmoothing.
-JDR simultaneously refine graph topology and node features by aligning the principal eigenspaces of their matrices, boosting GNNs across both homophilous and heterophilous networks.
-DropEdge randomly remove edges at each training epoch to act as both data augmentation and message-passing reduction, which is used to mitigate over-fitting and over-smooting problems.
+DropEdge randomly remove edges at each training epoch to act as both data augmentation and message‑passing reduction, which is used to mitigate over-fitting and over-smooting problems.
+FoSR is a preprocessing method, which aim to address the oversquashing issue by improving the graph connectivity. It adds edges by exploring the first order change in the spectral gap.
+BORF uses the Ollivier-Ricci curvature to rewire graph, where minimally curved edges causing the information bottlenecks should add connections and maximally curved edges leading to over-smoothing should be removed.
+SJLR combines the Jost–Liu Curvature of each edge with the embedding similarity between its incident nodes, and uses the weighted score as the probability for edge removal or addition.
+DHGR compares the neighborhood feature distribution and neighborhood label distribution between node pairs; edges connecting nodes with low similarity (heterophilous) are pruned, while edges between highly similar (homophilous) nodes are added.
+DHGR vs.
+THR.
+Although both methods essentially assess edge homophily or heterophily through feature and label differences, they follow distinct methodological lines.
+DHGR is a preprocessing approach that aggregates neighborhood features and derives local label distributions from pseudo-labels produced by a pre-trained model, a heuristic design without explicit theoretical grounding.
+By contrast, THR operates within the optimizing model, contrasting node representations and quantifying their homophily ratio disparity, thereby aligning with prior theoretical proofs and offering a more principled formulation.
 C.4 Hyperparameters
-In the subsection, we list the detailed hyperparameters used for the experiments and they are also provided in code. The hyperparameters can be found in Table 4.
-Table 4: Hyperparameters of TorqueGNN in 10 datasets.
-Datasets Lr Wd Dropout $L$ $t$ $r$ epochs
-Texas 0.001 5e-4 0.7 8 5 0.2 100
-Wisconsin 0.001 0.5 0.5 8 5 0.2 100
-Cornell 0.001 0.05 0.7 8 2 0.2 100
-Actor 0.001 0.05 0.1 8 5 0.2 100
-Citeseer 0.001 5e-8 0.1 8 2 0.2 100
-Cora 0.001 0.1 0.1 8 2 0.2 100
-Pubmed 0.001 0.08 0.1 12 5 0.2 100
-Tolokers 0.001 5e-8 0.1 2 2 0.2 500
-Questions 0.001 5e-8 0.1 2 2 0.2 500
-Penn94 0.001 5e-8 0.1 2 2 0.2 500
+In the subsection, we list the detailed hyperparameters used for the experiments and they are also provided in code. The hyperparameters can be found in Tables 6-7.
+Table 5: Hyperparameters of THR on GCN across 11 datasets.
+Datasets Lr Wd Dropout L t epochs Normalize Data Hidden Size
+Texas 0.05 0.0005 0.5 2 5 10000 Yes 32
+Wisconsin 0.05 0.0005 0.5 2 5 10000 Yes 32
+Cornell 0.05 0.0005 0.5 2 5 10000 Yes 32
+Actor 0.01 0.0005 0.5 2 2 10000 No 32
+Citeseer 0.01 0.0005 0.5 2 10 10000 Yes 32
+Cora 0.01 0.05 0.5 2 2 10000 No 32
+Pubmed 0.01 0.0005 0.5 2 2 10000 Yes 32
+Tolokers 0.005 5e-8 0.2 2 1 10000 No 32
+Questions 0.005 5e-8 0.2 5 1 10000 No 32
+Penn94 0.001 5e-8 0.5 2 1 10000 No 32
+Flickr 0.01 0.0005 0.5 2 1 10000 No 32
+Table 6: Hyperparameters of THR on GPRGNN across 11 datasets.
+Datasets Lr Wd Dropout L t epochs Normalize Data PPR Hidden Size
+Texas 0.05 0.0005 0.5 2 5 10000 Yes 1 32
+Wisconsin 0.05 0.0005 0.5 2 5 10000 Yes 1 32
+Cornell 0.05 0.0005 0.5 2 5 10000 Yes 0.9 32
+Actor 0.01 5e-8 0.5 2 2 10000 Yes 0.9 32
+Citeseer 0.01 0.0005 0.5 2 10 10000 Yes 0.1 32
+Cora 0.01 0.0005 0.5 2 5 10000 Yes 0.1 32
+Pubmed 0.05 0.0005 0.5 2 2 10000 Yes 0.2 32
+Tolokers 0.005 5e-8 0.5 2 1 10000 No 0.1 256
+Questions 0.05 5e-8 0.5 2 1 10000 No 0.1 32
+Penn94 0.01 0.0001 0.5 2 1 10000 No 0.1 32
+Flickr 0.01 0.0005 0.5 2 1 10000 No 0.1 32
+Table 7: Hyperparameters of THR on APPNP across 11 datasets.
+Datasets Lr Wd Dropout $\alpha$ $L$ $t$ epochs Normalize Data Hidden Size
+Texas 0.001 0.0005 0.7 0.05 8 5 100 No 512
+Wisconsin 0.001 0.5 0.5 0.05 4 5 100 No 512
+Cornell 0.001 0.05 0.7 0.5 8 2 100 No 512
+Actor 0.001 0.05 0.1 0.5 8 5 100 No 512
+Citeseer 0.001 0.05 0.4 0.5 8 10 100 No 512
+Cora 0.001 0.5 0.4 0.5 4 2 100 No 512
+Pubmed 0.001 5e-8 0.4 0.5 4 2 100 No 512
+Tolokers 0.001 5e-8 0.1 0.8 2 2 500 No 512
+Questions 0.001 5e-8 0.1 0.8 2 2 500 No 512
+Penn94 0.001 5e-8 0.1 0.8 2 2 500 No 512
+Flickr 0.001 0.5 0.1 0.8 2 2 500 No 512
 C.5 Experiments
-Ablation Study.
-To validate the effectiveness of our proposed torque metric, we conduct ablation studies comparing three edge-scoring schemes.
-As reported in Table 5, reliance on structural distance alone yields the poorest classification performance on most benchmarks, reflecting the fact that certain long-range connections convey meaningful semantics.
-Introducing the energy metric markedly improves accuracy, confirming its significance as a measure of edge perturbation strength.
-Moreover, on Tolokers and Pubmed, the energy-only variant achieves the highest accuracy, slightly surpassing the combined torque score, likely because some low-interference connections carrying critical semantic information are inadvertently pruned when the distance is too large.
-Crucially, our torque score, integrating both distance and energy, consistently outperforms either criterion besides Tolokers and Pubmed, thereby validating its superior ability to accurately identify disruptive edges, guide interference-aware rewiring, and enhance overall classification accuracy.
-Table 5: Ablation Study: Performance of TorqueGNN equipped with different edge-scoring metrics; the best results are shown in bold.
-Datasets Distance Energy Torque (Distance $\times$ Energy)
-Texas 88.61 (0.70) 89.16 (0.69) 91.11 (1.67)
-Wisconsin 79.20 (2.40) 84.60 (1.80) 89.60 (1.20)
-Cornell 81.94 (4.17) 83.05 (0.83) 89.44 (1.67)
-Actor 36.69 (0.40) 35.48 (0.07) 36.72 (0.16)
-Tolokers 72.46 (0.04) 76.11 (0.28) 75.28 (0.02)
-Penn94 81.11 (0.09) 81.20 (0.15) 81.30 (0.19)
-Questions 63.11 (1.26) 62.65 (0.53) 63.54 (0.85)
-Citeseer 73.35 (0.24) 76.53 (0.42) 76.84 (0.60)
-Cora 78.30 (0.59) 84.50 (0.00) 85.39 (0.30)
-Citeseer 73.35 (0.24) 76.53 (0.42) 76.84 (0.60)
-Pubmed 88.01 (0.11) 88.52 (0.16) 88.36 (0.08)
+Classification Results.
+Table 8 shows the performance gains brought by APPNP with diverse rewiring methods.
+We can observe that on most datasets, THR obtains the optimal performance, indicating its effectiveness.
+Table 8: Node classification results on benchmark datasets with APPNP as the backbone models: Mean ACC % (Standard Deviation %).
+The first- and second-best accuracies are highlighted in bold and underlined, respectively.
+Methods/Datasets APPNP FoSR BROF SJLR DHGR DropEdge-L THR
+Texas 49.17±3.30 78.04 (3.70) 73.53 (7.66) 81.57 (3.94) 78.04 (4.62) 72.75 (4.68) 84.12 (3.22)
+Wisconsin 47.60±4.54 74.26 (4.47) 75.00 (4.41) 83.53 (5.95) 73.82 (3.77) 71.91 (4.85) 87.79 (3.54)
+Actor 35.24 (0.56) 35.51 (1.42) 35.35 (1.28) 35.19 (1.13) 35.90 (1.16) 35.48 (1.12) 36.34 (0.87)
+Cornell 67.57 (5.54) 67.57 (5.54) 68.38 (7.46) 74.59 (5.16) 69.46 (8.80) 69.19 (7.27) 77.57 (8.02)
+Penn94 76.53 (0.28) 76.53 (0.28) OoM 79.73 (0.24) 79.10 (0.40) 76.13 (0.40) 82.56 (0.43)
+Flickr 57.26 (7.69) 56.31 (6.99) OoM 61.86 (5.17) 62.20 (1.10) 61.25 (3.30) 63.28 (1.56)
+Citeseer 74.02 (0.38) 77.65 (1.55) 77.65 (1.24) 77.25 (1.35) 76.89 (1.81) 77.69 (1.67) 78.74 (1.29)
+Cora 85.89 (1.19) 85.89 (1.19) 85.19 (1.87) 86.51 (1.59) 85.85 (1.79) 85.54 (1.14) 86.35 (1.61)
+Pubmed 87.19 (0.55) 87.19 (0.55) 87.16 (0.40) 88.84 (0.40) 88.47 (0.44) 87.66 (0.33) 88.31 (0.49)
+Tolokers 75.11 (0.74) 75.11 (0.74) OoM 78.46 (1.11) 75.33 (0.83) 74.64 (1.06) 79.29 (0.42)
 Parameter Sensitivity.
+Although $\alpha$ balancing the contribution of the learned high-order representation and the original input features originates from APPNP, THR modifies the graph structure over which propagation occurs.
+To examine how signal diffusion changes with respect to $\alpha$ under the rewired graph, we perform a sensitivity analysis shown in Figure 4, where a larger $\alpha$ increases the influence of the hidden representations.
+We observe that smaller heterophilous graphs (e.g., Texas and Actor), optimal accuracy is achieved at low $\alpha=0.05$, implying that raw node features provide sufficient discriminative power.
+In contrast, on larger or homophilous graphs, better performance is observed when $\alpha=0.5$, reflecting the necessity of high-order hidden representations to capture more complex community structures.
+Moreover, for all datasets, the best results are gained at a larger $\alpha$, which demonstrates the effectiveness of excavating deep features.
+Figure 4: Parameter sensitivity: Performance curves on five datasets with layers changing in $\{2,4,8,16,32\}$.
 Figure 5 explores the effect of network depth $L$.
-Across most datasets, performance peaks at $L=8$, thereby validating the effectiveness of our THR strategy for mining high-order information.
-Beyond eight layers, accuracy declines, primarily due to training challenges typical of deep networks, such as over-smoothing.
-While our rewiring mechanism mitigates over-smoothing to some extent, it is not explicitly engineered to resolve this phenomenon; overcoming depth-related bottlenecks remains a direction for future work.
-Figure 5: Parameter sensitivity: Performance curves on five datasets with layers changing in $2\}$ , 4 , 8 , 16 , 32 }.
-Appendix D Broader Impact Statement and Limitations
+For small graphs (Texas, Citeseer and Film), performance improves as the number of layers increases, since deeper networks are required to capture sufficient high-order information.
+In contrast, for large graphs (Tolokers and Flickr), the best performance is achieved with only two layers, indicating that shallow message passing already provides sufficiently discriminative representations.
+However, while APPNP can alleviate over-smoothing to some extent, it does not explicitly address this issue on these graphs; overcoming depth-related bottlenecks therefore remains an open direction for future research.
+Figure 5: Parameter sensitivity: Performance curves on five datasets with layers changing in $\{2,4,8,16,32\}$.
+Appendix D Broader Impact Statement
 This study aims to enhance message passing in graph neural networks through graph rewiring.
 As a result, it contributes to better performance and broader applicability of GNNs across a wide range of tasks, including recommendation systems, molecular property prediction, traffic forecasting, and social network.
-THR depends on sorting operations that impose considerable computational and memory burdens on large, densely connected graphs.
-Additionally, by assigning equal weight to distance and energy metrics, THR may inadvertently prune edges that carry essential semantic information and exhibit low interference when the distance is too large.
-How to optimally balance these two metrics in accordance with the characteristics of different datasets remains an open challenge.